### v1

# Evaluation of uncertainty estimations for Gaussian process regression based machine learning interatomic potentials

Abstract
Machine learning interatomic potentials (MLIPs) have seen significant advances as efficient replacement of expensive quantum chemical calculations for predicting molecular energies and forces. Uncertainty estimations for these machine learning models are crucial to quantify the additional model error they introduce and to leverage this information to find most informative training samples in active learning strategies. MLIPs that are based on Gaussian process regression provide predictive distributions with a standard deviation as a possible uncertainty measure. An alternative approach are ensemble-based uncertainties. Although these uncertainty measures have been applied to active learning, it has rarely been studied how well or whether they correlate with the actual error, and it is not always clear whether active learning strategies actually outperform randomized sampling strategies. In our study, we consider GPR models with Coulomb and Smooth Overlap of Atomic Positions (SOAP) representations as inputs to predict potential energy surfaces and excitation energies of molecules. We evaluate, how the GPR variance and ensemble-based uncertainties relate to the actual error and whether model performance improves by selecting the most uncertain samples from a fixed configuration space for training. For the ensemble based uncertainty estimations, we find that they often do not provide any information about the error. For the predictive GPR standard deviation, we find that in most cases predictions with an increasing standard deviation also have an increasing systematical bias, which is not captured by the uncertainty. In these cases, selecting training samples with the highest uncertainty leads to a model with a worse test error compared to selecting samples randomly. We conclude that confidence intervals for GPR-based MLIPs, which are derived from the predictive standard deviation, can be highly overconfident, as they do not capture the bias in the model predictions. Selecting samples with high GPR standard deviation leads to a model that overemphasizes the borders of the configuration space represented in the fixed dataset. This may result in worse performance in more densely sampled areas but better generalization for extrapolation tasks.
1 Introduction
Machine learning (ML) models that map atomic structures to corresponding properties have seen a significant increase in research interest over the past decade. These models promise to combine the accuracy of ab initio calculations with the speed, and scalability to large systems, of empirical potentials. Of particular interest are models that predict potential energies and possibly the associated forces, known as machine learning interatomic potentials (MLIPs). The idea is to perform highly accurate reference calculations, e.g. with density functional theory, for only a few molecular configurations. These samples then serve as training data for a machine learning model, which can predict new samples at substantially lower costs.
How well samples from different parts of the configuration space are predicted by an MLIP depends highly on its training data. By configuration space, we refer to the set of all possible spatial arrangements of atoms in an atomic system. Samples from areas of the configuration space that differ too much from those covered by the training data might not be predicted well. Because of this variability in prediction accuracy, it becomes desirable to provide an uncertainty estimation for every model prediction, which ideally gives a good estimation for the prediction error. A good error estimation could help to understand the range of a potential error of a prediction, and could be utilized in active learning (AL) strategies1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15; 16; 17. With AL, algorithms are referred to that aim to choose or create new training samples, which are maximally informative to the model. In the context of MLIPs, two important goals of AL exist. The first goal is to increase the data efficiency of the model training. Data efficiency refers to the reduction in the number of training samples needed to achieve comparable model accuracy within a fixed configuration space, compared to using randomly selected samples. A second goal is to sample the configuration space, without being fixed, in all areas which are relevant for the respective applications.
Many different ML approaches to design MLIPs exist. One class of MLIPs are message passing neural networks 18; 19; 20; 21 (MPNN), which contain features that are trained via message passing steps. In other approaches, instead of trained features, fixed representations are employed that either globally represent the whole system of atoms 22; 23; 24 or locally represent atomistic environments 25; 26; 27; 28. These representations provide important properties such as invariance against translation, rotation, and atom permutation29. Besides linear models 26; 30 and neural networks28; 31; 32, kernel-based models are commonly chosen as ML model 33; 22; 24. Within kernel-based models, many are variations of Gaussian process regression34 (GPR). GPR predicts the output for a new input by generating a Gaussian predictive distribution, where the mean represents the (deterministic) prediction and the variance provides a measure of uncertainty for the prediction. The GPR variance is commonly used as an uncertainty measure in AL schemes for GPR-based MLIPs9; 16; 3; 1. Besides the GPR variance, other uncertainty estimations are conceivable for GPR. A common approach in ML is to construct ensembles of slightly different models and use the variance of their predictions as an uncertainty measure. Model ensembles can be based on different initializations of weights 35 or they can be generated by training on different subsets of the training data36. For neural networks, using different weight initializations is reasonable because these models are parametric and sensitive to their initial conditions. In contrast, for GPR, it is more reasonable to use different subsets of the training data because GPR is a non-parametric method that relies directly on the training data to make predictions. Uteva et al.9 trained two models on half of the initial training set each. Similarly, a bootstrap aggregation36; 37 ensemble, which is commonly used in ML, would be possible for GPR. Here multiple models are trained on different subsets of the training data, which are generated by randomly sampling indices from the original dataset. Our focus will be on GPR-based MLIPs with GPR variance and ensemble approaches as uncertainty measure or predictive distributions.
GPR models, as well as GPR ensembles, provide predictive distributions with a mean and a standard deviation. The mean of this distribution may serve as point prediction, which is the model’s actual predicted value. The standard deviation is a measure of variability or uncertainty in the prediction. An important question is, to which extend such uncertainty measures are correlated to the prediction error of a model, thus whether they would serve as error estimator. This requires a careful evaluation. Evaluating uncertainty estimations of MLIPs is a challenging problem with a large variety of approached being reported in the literature. Tran et al.38 distinguished between three aspects of uncertainty estimation: calibration, sharpness and dispersion. Calibration describes if the uncertainty estimation is correct, in the sense that the true error values follow the predicted distribution. Sharpness characterizes the overall confidence of a model’s predictions, and generally, a model with more confident predictions is preferred. Dispersion, on the other hand, refers to the spread of the uncertainty estimates. A model with high dispersion shows a wide variability in confidence across different predictions, allowing it to differentiate between samples predicted with high accuracy and those potentially having high errors. Tran et al.38 were analyzing these aspects with a variety of metrics. In other studies, often the uncertainty estimations are plotted against the actual or absolute errors3; 39; 40; 41; 10, which, in principal, gives an intuitively accessible picture off all three aspects. Kellner et al. 39 constructed an example with synthetic data to show how an ideal uncertainty versus error double-logarithmic plot would look like and illustrated various cases of incorrect uncertainty estimations. They then analyzed, if uncertainty versus error plots for non-synthetic data show the correct shape. Busk et al. 40 sorted all predictions in equally sized bins and calculated metrics like the z-score variance and the RMSE for all samples of one bin. In this work, we use binning to analyze the error distributions of predictions with certain uncertainties.
In the literature, uncertainty estimations of MLIPs are applied in a large variety of different active learning schemes. Many of these create new configurations without explicitly restricting the configuration space. A particularly important AL scheme is on-the-fly AL. Here, an uncertainty is calculated at every point during an MD run with the MLIP to recognize if the MLIP runs into configuration areas which are poorly sampled in the training data and for which the MD run might become instable1; 2; 3; 4; 5; 6; 7. Other schemes are optimizing the uncertainty estimation multiplied by a weight function to create new training configurations15; 16; 17. However, when using uncertainty estimation to guide the exploration of a flexible configuration space, it becomes challenging to assess the actual impact of active learning and the chosen uncertainty metric in comparison to generating samples randomly. Therefore, we restrict ourselves to a simple, very common AL approach from standard ML literature, which samples a fixed configuration space. This approach is often referred to as uncertainty sampling 42; 9. Uncertainty sampling can be applied in setups, where a large amount of unlabeled data exists and generating the respective labels is expensive. For a small random subset of the data, labels are generated. Then this data is used to train an initial model. The uncertainty estimation of this model is used to iteratively select samples from the remaining pool of unlabeled samples for which the labels are generated and which are then added as additional training samples. In the general ML literature, uncertainty sampling is often applied to classification problems, but it is also used for regression problems 43. It was shown that uncertainty sampling can increase the data efficiency by a multiple, although the effect strongly depends on the dataset and the uncertainty estimation that is used. In this study, uncertainty sampling is applied to analyze the benefit of the discussed different uncertainty estimation methods on AL.
In literature, uncertainty estimations of GPR based MLIPs are often evaluated by directly examining the performance when applied to AL without explicitly analyzing how the uncertainty estimations correlate with the actual error9; 1; 16. These studies give a mixed picture on the effect of AL on training GPR-based MLIPs and it is often not clear, whether AL strategies provide an actual improvement over random sampling. The correlation between uncertainty estimations of GPR-based MLIPs and the actual error has rarely been studied, explicitly. Vandermause et al.3 plotted their Gaussian process uncertainty against the true absolute error and reported that most true absolute errors fall within twice the Gaussian process uncertainty. They further showed how their active learning method, based on the Gaussian process uncertainty, provides actual improvement over random sampling. Bartok et al. 44 plotted uncertainty estimations against the true absolute error for a dataset of Ar dimers and trimers, finding that maximizing the leave-one-out cross-validation likelihood for hyperparameter optimization can be advantageous for uncertainty estimation, compared to maximizing the marginal likelihood or manual hyperparameter tuning. Tran et al. 38 evaluated different aspects of uncertainty estimations for an exemplary dataset using, among others, a GPR-based MLIP. Various GPR-based MLIPs exist, which differ in how their kernel is constructed and which representations are used. Therefore, also the GPR variance differs and it is not clear how the GPR variances of different models compare as uncertainty estimations. Furthermore, even if an uncertainty estimation shows a meaningful correlation with the errors, translating this relationship into enhanced data efficiency for AL runs remains uncertain.
The objective of this study is hence to evaluate uncertainty measures for GPR-based MLIPs. We consider a GPR model with global Coulomb representations22 as inputs as well as an atomistic GPR model with SOAP25 representations. We test the prediction of ground state energies from molecules of well-established benchmarks, allowing for easy comparison and categorization of our results. We further predict excited state energies for a trajectory of porphyrin in gas phase as a large and complex molecule with a potentially complex configuration space. To evaluate the uncertainty estimations, we test how the GPR variance, as well as ensemble uncertainties, of different GPR-based MLIPs, correlate with the actual error for unseen samples. This is done by analyzing the distributions of errors of predictions with different uncertainties. It is further investigated how different uncertainty measures perform, when applied to AL. Uncertainty sampling is chosen as AL algorithm for the reasons discussed above. To understand, how much information could be maximally extracted from the candidate pool, when using uncertainty sampling, we always also provide a reference AL experiment. In that reference AL experiment, exact information on the error is used, in contrast to using an uncertainty measure.
This work is structured as follows. In Section 2, Gaussian Process Regression (GPR) is introduced, and the use of Coulomb and SOAP representations to construct GPR-based MLIPs is described. The different uncertainty estimations that are applied are described and we give a detailed description of how we evaluate them. Further details are given on how uncertainty sampling was applied and which data is used. In Section 3 we analyze the correlation of uncertainty estimations and errors for the different datasets. It is further investigated how the observed correlations transfer when applied to uncertainty sampling. An overall discussion of the results is given in Section 4. Conclusions with an outlook are given in Section 5.
2 Methods
2.1 Gaussian process regression
In GPR34, a prior probability distribution over functions is defined using a Gaussian process. We want to find a posterior predictive distribution from training inputs $\mathbf{x}_{i}$ and corresponding output measurements $y_{i}$, with which we can predict outputs for unknown inputs x. Our output measurements will always be noisy. This can be considered by assuming that the actual function values differ from the measurements with
$y_{i}=f\left(\mathbf{x}_{i}\right)+\epsilon_{i},\qquad\text{with}\,\epsilon_{i}\sim\mathcal{N}\left(0,\sigma_{n}^{2}\right)\,,$ (1)
where $\epsilon_{i}$ are independent, identically distributed random variables drawn from a Gaussian distribution with zero mean and variance $\sigma_{n}^{2}$. The GPR framework provides a predictive distribution for the output corresponding to any input x. The point prediction for the output of an input $\mathbf{x}$ is the mean of this predictive distribution, given as
$m\left(\mathbf{x}\right)=\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{y}\,.$ (2)
The properties of this function are determined by the covariance function $k\left(\mathbf{x},\mathbf{x^{\prime}}\right)$, which measures the similarity between two inputs $\mathbf{x}$ and $\mathbf{x^{\prime}}$. $\mathbf{X}$ is a matrix with the training inputs as rows, $\mathbf{y}$ is a vector with the corresponding outputs. The vector $\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)$ contains all pairs of kernel evaluations between the input $\mathbf{x}$ and the training inputs $\mathbf{X}$, whereas $\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)$ is the matrix of all pairwise kernel evaluations on the training inputs $\mathbf{X}$.
In our case, an input x to the model can be either a global representation of the atomistic system or local representations of the atomistic environments. The corresponding output $y$ is a scalar energy value. For our tests, we chose a Gaussian covariance function
$k\left(\mathbf{x},\mathbf{x^{\prime}}\right)=\sigma_{f}^{2}\text{exp}\left(-\frac{1}{2}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)^{T}\mathbf{M}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)\right)\,,$ (3)
which is one of the most common choices. We used a different length-scale for every feature, known as automatic relevance determination34, to increase the accuracy and flexibility of the model. Here $\mathbf{M}$ is a diagonal matrix $\text{diag}\left(\frac{1}{l_{1}^{2}},...,\frac{1}{l_{D}^{2}}\right)$, with $D$ the number of features, and $\sigma_{f}^{2}$ is the output-scale. We optimized the hyperparameters $\sigma_{f}^{2}$ and $\sigma_{n}^{2}$, as well as length-scales inside the covariance function via maximum marginal log-likelihood, details can be found in Section S2. For all our tests, we used the GPR implementation of GPyTorch 45.
In the case of a global representation as inputs $\mathbf{x}$, the covariance function of Eq. 3 is directly applied to these inputs. For our study, we incorporate Coulomb matrices22 as a global representation. For Coulomb matrices, to ensure invariance against atom indexing, often a sorting is carried out, which results in discontinuities, which may cause challenges in the construction of ML models 29. In our tests, we ignored invariance against atom permutation and omitted sorting because, for all tasks, a system with fixed atoms given in a fixed order was given.
As a local representation, we used Smooth overlap of atomic positions (SOAP) 25. It is commonly used and GPR models with SOAP representations as inputs have shown high accuracy 29. For atomistic representations, an input x will consist of representations $\textbf{x}_{j}$ for every atomistic environment of the system. The covariance function between two inputs x and x’ is built by summing over an atomistic covariance function
$k\left(\mathbf{x},\mathbf{x^{\prime}}\right)=\sum_{j,l=1}^{n_{\text{atoms}}}k_{\text{atomistic}}\left(\textbf{x}_{j},\textbf{x}^{\prime}_{l}\right).$ (4)
[…]
2.2 Uncertainty estimations
As discussed in the introduction, GPR and ensembles of GPR models provide a probabilistic predictive distribution. In general, we will describe this distribution via a pair
$\mathcal{D}(m(\mathbf{x}),u(\mathbf{x})^{2})\,,$ (5)
with a given mean $m(\mathbf{x})$, and a standard deviation $u(\mathbf{x})$, which we consider as uncertainty measure in here. Note that we may interpret the pair from eq. 5 as the parameters describing a Gaussian distribution. In fact, the predictive distribution of a GPR model is anyway Gaussian. Moreover, the predictive distributions of GPR ensembles are often assumed to be Gaussian too40; 47, which is, according to the central limit theorem, at least true for a large number of ensemble models47.
For the case of the standard GPR model, the predictive distribution is, using notation from eq. (5), given by a pair
$\mathcal{D}_{G}(m_{G}(\mathbf{x}),u_{G}(\mathbf{x})^{2})\,,$ (6)
where $m_{g}(\mathbf{x})$ is the GPR mean from eq. 2 and we have
$u_{g}\left(\textbf{x}\right)=\sigma\left(\textbf{x}\right)$ (7)
for all $\mathbf{x}$, with the GPR variance
$\sigma^{2}\left(\mathbf{x}\right)=k\left(\mathbf{x},\mathbf{x}\right)-\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{K}\left(\mathbf{X},\mathbf{x}\right)\,.$ (8)
Note that the GPR variance is only calculated from the kernel function $k(\cdot,\cdot)$, the inputs $\mathbf{X}$ of the training set, i.e. the sample locations in the feature space, and the noise variance $\sigma_{n}^{2}$. Hence it does not depend on the training outputs, i.e. the trained energies. Intuitively, the GPR variance is high, when $\mathbf{x}$ is far from the training inputs, as measured by the GPR model. In the variance, see eq. 8, GPR already accounts for the noise $\sigma_{n}^{2}$ in the observations, therefore, we do not add this noise to our uncertainty estimation, like it was done by Vandermause et al.3.
The two set approach, proposed by Uteva et al.9, was motivated by the principles underlying the jackknife and bootstrap methods36. We use for its predictive distribution the notation
$\mathcal{D}_{2}(m_{2}(\mathbf{x}),u_{2}(\mathbf{x})^{2})\,.$ (9)
This distribution is obtained by splitting the initial training set into two sets which are used to train two GPR models with predictions $m^{\left(1\right)}\left(\textbf{x}\right)$ and $m^{\left(2\right)}\left(\textbf{x}\right)$. The mean over both models then becomes the predictive mean
$m_{2}(\mathbf{x})=\frac{1}{2}(m^{\left(1\right)}\left(\mathbf{x}\right)+m^{\left(2\right)}\left(\mathbf{x}\right))\,,$ (10)
while the difference in the predictions of the two models
$u_{2}\left(\textbf{x}\right)=\left|m^{\left(1\right)}\left(\textbf{x}\right)-m^{\left(2\right)}\left(\textbf{x}\right)\right|\,,$ (11)
is the uncertainty estimation. The square of this particular uncertainty measure can be interpreted as a very rough estimate of the model’s variance with respect to changes in the training set, including the energies. In other words, it very roughly approximates the variance in the bias-variance decomposition of ML models, which is common in ML literature48.
An extension of this is the bootstrap aggregation ensemble36, which leads to the predictive distribution
$\mathcal{D}_{b}=(m_{b}(\mathbf{x}),u_{b}(\mathbf{x})^{2})\,.$ (12)
It is commonly used for neural networks. To calculate the predictive distribution, multiple subsets of the training data are generated by randomly sampling indices from the original dataset, allowing for duplicates. Then, the subsets are used to train an ensemble of multiple models $m^{\left(l\right)}\left(\textbf{x}\right)$. The mean is then
${m}_{b}\left(\textbf{x}\right)=\sum_{l=1}^{N}m^{\left(l\right)}\left(\textbf{x}\right)\,,$ (13)
and the uncertainty is calculated by the standard deviation of their different predictions $m^{(l)}\left(\textbf{x}\right)$,
${u}_{b}\left(\textbf{x}\right)=\sqrt{\frac{1}{N-1}\sum_{l=1}^{N-1}\left(m^{\left(l\right)}\left(\textbf{x}\right)-{m}_{b}\left(\textbf{x}\right)\right)^{2}}\,,$ (14)
where $N$ is the number of models. Note that the variance of the bootstrap ensemble is an even better approximation of the variance of the bias-variance decomposition.
As a slight change towards the classical definition of the two-sets and bootstrap approach, we do not use the ensemble mean as mean for their predictive distributions, since we found whether we use the mean of the ”full” GPR model or the ensemble mean does not make a difference for the essential results on the correlation between error and uncertainty estimation. A comparison can be found in Section S2.
2.3 Evaluating uncertainty estimations
The main objective of this study is to analyze the relationship of uncertainty estimations, such as GPR variance, and the actual prediction error. In particular, it is of interest to understand whether uncertainty estimations of ML models can serve as error estimator. A straight forward way to analyze this is to calculate uncertainty estimations and errors of a set of $n_{\text{test}}$ unseen test configurations and plot the uncertainty estimations against the errors.
To better understand, what we could potentially expect from uncertainty estimations in an ideal setting, let us assume we would have an approach that would construct an optimal predictive distribution
$\mathcal{D}_{opt}(m_{opt}(\mathbf{x}),u_{opt}(\mathbf{x})^{2})\,,$
where the true values $y(\mathbf{x})$ related to predictions $m_{opt}(\mathbf{x})$ are actually following the distribution
$y(\mathbf{x})\sim\mathcal{N}(m_{opt}(\mathbf{x}),u_{opt}(\mathbf{x})^{2})\,.$
It is then quite immediate that, under this thought experiment, the prediction error $\epsilon(\mathbf{x})$ of the model for all test samples would behave like
$\epsilon(\mathbf{x})=y(\mathbf{x})-m_{opt}(\mathbf{x})\sim\mathcal{N}(0,u_{opt}(\mathbf{x})^{2})\,.$
In other words, the error of the model would be normally distributed with a mean of 00 and a standard deviation of $u_{opt}(\mathbf{x})$. In principle, this what we assume, if we consider the standard deviation of the predictive distribution to match, or to be perfectly correlated, to the error of the model. Consequently, our subsequent evaluation of the uncertainty estimations aims at testing for a given uncertainty measure $u(\mathbf{x})$, how closely the prediction errors $\epsilon$ would fulfill
$\epsilon\sim\mathcal{N}(0,u(\mathbf{x})^{2}))\,.$ (15)
If the observed error distribution closely aligns with eq. (15), valid confidence intervals, which provide a range that is expected to contain the true error with a specified probability, can be derived from the uncertainties of predictions. To validate if the errors of predictions with the same predictive variance are distributed according to eq. (15), we can separate all predictions by their uncertainty values into $n_{\text{bins}}$ equidistant bins with intervals
$I_{\beta}=\left[\left(\beta-1\right)\Delta u,\,\beta\Delta u\right),\qquad\text{with}\qquad\beta=1,...,n_{\text{bins}},$ (16)
where $\Delta u$ is the bin width. We then assume that all predictions inside one bin approximately have the center of that bin
$u_{\beta}(\mathbf{x})=\left(\beta-\frac{1}{2}\right)\Delta u$ (17)
as same fixed uncertainty. To evaluate whether the true error distributions align with the model’s prediction, one can calculate the bin mean and bin standard deviation of the errors $\epsilon_{\beta}$ within each bin. The mean of $\epsilon_{\beta}$ should be zero and the standard deviation should be $u_{\beta}\left(\textbf{x}\right)$, according to eq. (15). If the bin mean and bin standard deviation align with the model prediction, and the errors are also distributed according to a Gaussian, as it is predicted, also the confidence intervals, which are derived from the predictive distributions would be correct. We will refer to plots in which we plot the uncertainty estimation against the errors and mark the bin means and standard deviations of errors as correlation plots.
To illustrate a model, which perfectly aligns with eq. (15), Fig. 1 shows a correlation plot for data, where the targets were drawn from the predictive distribution of the GPR model. The distribution of errors can differ from this case, and therefore from Eq. 15, in two ways: the calculated mean of the errors is not zero and the calculated standard deviation of the errors is not $u_{\beta}\left(\textbf{x}\right)$. If the error mean differs from zero this indicates that the predictive distribution is shifted with a constant bias compared to the true distribution of targets. If the error standard deviation is smaller or larger than $u_{\beta}\left(\textbf{x}\right)$ it indicates, that the predictive distribution is either to narrow or broad, respectively.
Figure 1: Correlation plot for sinus data with synthetic GPR test targets. To generate the data a GPR model was trained with noisy data points from a sinus function. Test targets where drawn from the predictive distributions of the GPR model. Thereby data is generated for which the GPR model perfectly captures the underlying data-generating process. For all test samples the predicted GPR standard deviation is plotted against the actual error of the prediction. All predictions are separated by their standard deviation into equidistant bins. Mean and standard deviation of error distributions are shown for every bin.
2.4 Uncertainty sampling
With correlation plots, we can get a picture of different aspects that determine the quality of uncertainty estimations. As an additional study, we utilize the uncertainty estimations in an AL scheme, namely uncertainty sampling 42 to see, how effective is these estimations are in practical use. In uncertainty sampling, a small random subset of data is selected, and labels are generated for this subset. In our case, these labels are obtained via quantum chemical calculations of potential energies. This labeled data is then used to train an initial model. The uncertainty estimation of this model is used to iteratively select samples from the remaining pool of unlabeled samples for which the labels are generated and which are then added as additional training samples. See Algorithm 1 for pseudocode.
To analyze, how much uncertainty sampling with different uncertainty estimators increases data efficiency, we kept a test set separate from the data pool. In every iteration, we calculated the mean absolute error (MAE) for the test samples. Additionally, in every iteration, we calculated the maximum absolute error and the variance of the absolute errors of the test samples, to understand how consistently the model is performing across different data points. We compare adding samples via uncertainty sampling with adding samples randomly. We expect the highest increase in data efficiency, if we would have an uncertainty estimation that estimates the true error correctly. Therefore, to understand how much data efficiency could be achieved at maximum, we also iteratively added the sample with the largest absolute error in each iteration.
If, instead of adding one sample per iteration, batches of samples are added, there is a risk of adding multiple very similar samples, for all of whom the model has a high uncertainty. Although methods exist that balance informativeness, diversity, and representativeness11; 12, we do not use batch active learning approaches. Our goal to isolate the influence of the uncertainty estimation and get an accurate picture of how much additional information is extracted with different uncertainty estimations.
Algorithm 1 Uncertainty Sampling
0:  Unlabeled data pool $U$, initial sample size $n_{\text{init}}$, number of iterations $n_{\text{iter}}$
1:  Randomly select $n_{\text{init}}$ samples from $U$ and obtain their labels to form initial labeled dataset $L$
2:  $U$ = $U$$L$
3:  Train initial model $m$ using labeled data $L$
4:  for $i=1$ to $n_{\text{iter}}$ do
5:     Compute uncertainty for each sample in $U$ using model $m$
6:     Select sample $x^{*}$ from $U$ with highest uncertainty
7:     Obtain label $y^{*}$ for selected sample $x^{*}$
8:     $L=L\,\cup\,(x^{*},y^{*})$
9:     $U=U$$x^{*}$
10:     Retrain model $m$ using updated labeled dataset $L$
11:  end for
12:  return Trained model $m$
2.5 Data
MD1749 is a benchmark dataset of long molecular dynamic trajectories, calculated with density functial theory (DFT), for 10 small molecules. The revised MD17 (rMD17) dataset50 recalculated the energies as the original values were found to be noisy. Additionally the total number of snapshots from the separate trajectory was reduced to a selection of less correlated snapshots. We selected the rMD17 dataset for our tests because it is a well-established benchmark, allowing for easy comparison and categorization of our results. We performed tests for potential energies of Benzene, as a highly symmetric molecule and Aspirin, as a more complex molecule.
The WS22 dataset51 consists of distributions of configurations obtained via Wigner sampling for 10 different organic molecules, which aims to cover an extensive configuration space. This is important for our tests to understand how uncertainty estimations can distinguish between areas of configuration spaces that are predicted accurately and those that are predicted poorly across a wide range of configurations. We performed tests for potential energies of SMA and O-HBDI.
Additionally, we considered porphyrin as a large and complex molecule with a potentially complex configuration space. The dataset consists of 100,000 snapshots of a quantum molecular dynamic trajectory calculated with Density Functional Tight Binding (DFTB). To calculate the DFTB trajectory we took the initial structure from ChemSpider52 (ID: 4086). After DFT (B3LYP/6-31G) geometry optimization, 150,000 fs ground state dynamics at 300 K were performed with the DFTB+ package53, and 100,000 frames were extracted from the end of the trajectory with a stride of 1 fs. We performed our tests for the corresponding excitation energies which we calculated with time-dependent long-range corrected DFTB (TD-LC-DFTB).
Figure 2: Chemical structures of the molecules used in the tests.
3 Results
For our experiments, we randomly select 1000 training samples and 2000 test samples from every dataset, and the remaining samples form a candidate pool. For every dataset, we use the 1000 training samples to optimize the hyperparameters. To generate the correlation plots, we calculate the uncertainty estimations and prediction errors of the respective GPR model for the candidate pool samples and plot them against each other. We separated the samples into equidistant bins by their uncertainty estimations and calculated the bin means and standard deviations of the errors. For the different datasets, we manually selected the bin width based on the range of predicted uncertainties. The exact bin widths can be found in Section S3. We compare the calculated bin means and standard deviation with the theoretical values of Eq. 15. We only show the bin mean and standard deviation, if the number of samples inside one bin is greater than 50. For better visibility, we manually adjust the axis limits of the correlation plots, therefore a few scatter points are not displayed. For all uncertainty sampling tests, we use the fixed optimized hyperparameters and the test set, as described above. From all samples except the test samples we randomly choose 200 samples as initial training samples.
3.1 rMD17 benzene
3.1.1 Correlation between uncertainty estimate and prediction error
(a) GPR with Coulomb
(b) GPR with SOAP
Figure 3: Correlation of uncertainty and error across different uncertainty estimations for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset. The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq. 15.
Fig. 3 shows correlation plots of different uncertainty estimators for benzene of the rMD17 dataset. In Fig. 3(a), results for GPR with Coulomb representations are presented. For the GPR standard deviation and the bootstrap uncertainty, an increasing uncertainty correlates with a broader distribution of errors, as seen by the increasing bin standard deviations. Additionally, the bin means deviate further from zero with increasing uncertainty, indicating a bias in the model’s predictions. This suggests that both uncertainty estimators can identify samples for which the model has a high bias. The correlation between higher uncertainty and more broadly distributed errors as well as a higher bias means that for predictions with higher uncertainty higher errors are more likely. Still the observed error distributions do not align with what is predicted by the GPR models, making error ranges derived from these uncertainty estimations inaccurate. The systematical bias can not be capture by the uncertainty estimations. Besides that, for the GPR standard deviation, the bin standard deviations are smaller then the theoretical values, with an increasing deviation with increasing GPR standard deviation. Therefore, the errors are more narrowly distributed than predicted, which may partly compensate for the uncaptured bias when calculating confidence intervals. For the bootstrap uncertainty, for bins with lower uncertainty, the bin standard deviations exceed the theoretical values. Therefore the errors are distributed more broadly than predicted which will result in underconfident error intervals. For bins with higher uncertainty, the bin standard deviation aligns with the theoretical values, while the bin means deviate from zero, which will result in overconfident confidence intervals. Meanwhile, the two sets uncertainty shows errors consistently distributed around zero across all bins, with nearly identical standard deviations, providing no meaningful information about the error.
In Fig. 3(b), results for GPR with SOAP representations are shown. Again for the GPR standard deviation and the bootstrap uncertainty, an increasing uncertainty correlates with broader distributions of errors, therefore higher errors become more likely with increasing uncertainty. In contrast to GPR with Coulomb matrices, the bin means for both uncertainty estimators are close to zero and do not systematically increase with increasing uncertainty. Therefore, the GPR model has almost no systematical bias in the prediction. For the GPR standard deviation also the bin standard deviations are close to the theoretical value, therefore confidence intervals for predictions derived of the uncertainty would be rather accurate. For the bootstrap aggregation, the predicted error distributions for low uncertainties are slightly too broad, while error predictions of samples with higher uncertainties become too narrow with increasing discrepancy to the theoretical value for increasing uncertainty. This results in a less significant correlation between the uncertainty and a high likelihood for high errors than observed with the GPR standard deviation. Similar as for Coulomb matrices, the error distributions for the bins of the two set uncertainty do not differ for different uncertainties. Therefore, the two sets uncertainty again does not provide any meaningful information about the error.
3.1.2 Active learning via uncertainty sampling
Fig. 4 shows uncertainty sampling results for benzene of the rMD17 dataset. In Fig. 4(a), results for GPR with Coulomb representations are shown and in Fig. 4(b) results for GPR with SOAP representations are shown. Adding samples based on maximum uncertainty estimations does not lead to any improvement in MAE, and in most cases, it also fails to improve the maximum absolute error and test variance. Only in the range of 900 to 1200 training samples for GPR with SOAP, the model with randomly added samples shows worse test variance and maximum absolute error compared to the model where samples with the highest uncertainty estimations were added. The correlation analysis shows how some of the uncertainty estimators provide relevant information about the error. E.g. when the GPR standard deviation is used as uncertainty estimator, a higher uncertainty correlates with a higher likelihood for larger errors. For GPR with SOAP, the observed error distributions actually align well with what is predicted with the GPR standard deviation. For GPR with Coulomb, there is a correlation between predictions with the highest GPR standard deviations and predictions that have the highest bias. Still, neither for GPR with Coulomb nor for GPR with SOAP uncertainty sampling with GPR standard deviation does result in higher data efficiency. In both cases, it can be seen that adding samples with the maximum absolute error improves all metrics significantly, compared to adding samples randomly. This shows that, in theory, there is a significant potential for a higher data efficiency. With the absolute error, or a potential uncertainty estimator which linearly correlates with the error, a prediction with a high error has a high uncertainty. In contrast, if the uncertainty estimator is the standard deviation of a predictive distribution, the errors of predictions with a high uncertainty follow a distribution in which high errors have a high likelihood but still also low errors are likely. This means many samples identified as maximally uncertain actually have a low error. By adding these samples, the model becomes certain about them, but it was predicting them with a low error already before. This might explain why no improved data efficiency is observed with the GPR standard deviation or bootstrap uncertainty as uncertainty estimator.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure 4: Uncertainty sampling results for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally, we randomly added a new sample in every iteration, and we added the sample with the highest actual absolute error. During each iteration, we compute the MAE, the maximum absolute error and the variance for the test set.
3.2 Comparison for several molecules
3.2.1 Correlation between uncertainty estimate and prediction error
In Figure 5 correlation plots for the GPR standard deviation for different datasets are shown. In Figure 5(a), results for GPR with Coulomb are presented. In all cases, the bin means of bins with higher uncertainties deviate from zero, indicating a systematical bias of the respective predictions. Like it was observed for benzene, there is a correlation between increasing GPR standard deviation and increasing bias. In most cases, the bin standard deviation aligns rather well with the theoretical values. However, for SMA, it is observed that the bin standard deviations for bins with higher GPR standard deviations are smaller than the theoretical values. The deviation increases with increasing GPR standard deviation. The same can be observed to a smaller extent for O-HBDI. Therefore, the errors are more narrowly distributed than predicted. This may partly compensate for the uncaptured bias when calculating confidence intervals. Still, it can be observed that for predictions with high GPR standard deviation, the high bias dominates, which would lead to overconfident confidence intervals.
In Fig. 5(b), results for GPR with SOAP are presented. In all datasets, most of the bin means differ from zero, indicating that predictions with the GPR standard deviation of such bins have a bias. For SMA and O-HBDI, an increasing bias correlates with increasing uncertainty. For Aspirin, a clear correlation between increasing uncertainties and increasing bias cannot be observed. For SMA and Aspirin, the bin standard deviations align rather well with the theoretical values. Still, because of the uncaptured bias, resulting confidence intervals would be overconfident. For O-HBDI, the bin standard deviations for bins of low uncertainty align with the theoretical values, while for bins of high uncertainty the error distribution is much broader than predicted by the GPR model. This, combined with the uncaptured bias, would lead to highly overconfident confidence intervals.
Overall, the correlation between higher uncertainties and more broadly distributed errors, as well as a higher bias in most cases, show that there is a clear correlation between uncertainty and a higher likelihood for higher errors. Still the bin means and bin standard deviations of the error distributions do often not align with the theoretical values which are predicted by the GPR model, therefore no accurate confidence intervals can be derived from the GPR standard deviation. Specifically, predictions with a high GPR standard deviation often have a high bias, which dominates the error and is not captured by the GPR standard deviation, which would lead to highly overconfident confidence intervals for such predictions.
Here, we only compare correlation plots for the GPR standard deviation across different datasets, as it is the most widely used and discussed uncertainty estimator in the context of GPR based MLIPs. Results for two sets and bootstrap aggregation uncertainty estimations can be found in the supplementary material Section S4. In all cases, for the two sets uncertainty, the same picture as for benzene was observed: in all bins, the errors are distributed around zero with almost the same bin standard deviation. Therefore, again, the two sets uncertainty does not provide any meaningful information about the actual error. Similarly, correlation plots for the bootstrap aggregation uncertainty for Aspirin and SMA show very similar error distributions for all uncertainties, although a slight correlation between uncertainty and how much the bin means differ from zero, as well as how broad the error distributions are, can be observed. For O-HBDI, the bin standard deviations are larger than the theoretical values, with a correlation between increasing deviation and increasing uncertainties. The bin means of bins with large uncertainties also differ from zero.
3.2.2 Active learning via uncertainty sampling
In Fig. 6, the MAE for every iteration of uncertainty sampling for different datasets is shown. For all cases, beside porphyrin with Coulomb matrices, it can be observed that uncertainty sampling with the GPR standard deviation results in a worse MAE than adding samples randomly. Among all cases, porphyrin with Coulomb matrices is the only one where uncertainty sampling results in a similar MAE than adding samples randomly. In contrast to the other cases, similar to benzene with SOAP, porphyrin showed an error distribution in the correlation plot that aligns well with the theory. The two sets uncertainty performs very similar to adding samples randomly, what was expected as the correlation plots show how they do not provide any meaningful information about the error. Uncertainty sampling with the bootstrap aggregation uncertainty leads to a worse MAE compared to random sampling in all cases beside SMA with SOAP. Results for the maximum absolute test error and the test variance can be found in Section S5. The maximum absolute error of the test set improves for uncertainty sampling with GPR standard deviation in all cases, besides for aspirin. The test variance in most cases is similar to random sampling or deteriorates, only for O-HBDI a slight improvement can be observed. When training samples with the highest absolute error were added, all three metrics improved significantly for all cases, showing that in principle there is quite some potential for an improved data efficiency.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure 5: Correlation of the GPR standard deviation and error across different datasets for GPR with Coulomb and GPR with SOAP. The samples were separated into equidistant bins by their uncertainty, and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq. 15.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure 6: Uncertainty sampling results for GPR with Coulomb and SOAP representations for different datasets. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error. During each iteration, we compute the MAE.
3.3 Correlation analysis after applying active learning
It may be surprising that, in most cases, uncertainty sampling with the GPR standard deviation leads to a significantly worse MAE on the test set compared to random sampling. In Figure 5, we observe that predictions with the highest GPR standard deviation are having the highest bias. This would suggest that adding these configurations as training samples should be beneficial. The most extreme case is GPR with Coulomb matrices for aspirin, where the MAE after 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty is worse than for the initial model. To understand this behavior we show in Fig. 7 the correlation plots for Aspirin of a model trained via 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty estimation, as well as after iteratively adding 800 training configurations with the highest absolute error and adding 800 additional training configurations randomly. Similar to the model trained with 1000 random samples of Fig. 5, the initial model, which is trained with 200 random training configurations, shows an increasing negative bias for predictions with increasing GPR standard deviation, after GPR standard deviation exceeds a certain threshold. For predictions below this threshold, there is a range of rather low GPR standard deviation where the predictions have a positive bias. After adding 800 configurations randomly, this general form is still present. After uncertainty sampling, on the other side, the bias of predictions with a high standard deviation vanishes, while the positive bias, as well as the standard deviation of the errors, of predictions with low standard deviation, increases. A possible explanation for this is that, if we add the configuration for which the model predicts the highest GPR standard deviation in every iteration, we add the configuration which is most distant to the current training set. After many iterations, we have a training set with configurations which are maximally distant to each other. The training set now covers the borders of the configuration space, for which the initial model was showing a high bias, very well. However, predictions for some areas of the configuration space, which are less distant to the initial training configurations, are getting even worse, compared to the initial model with much less training configurations, when the training set overemphasizes the edges. In contrast to training via uncertainty sampling and adding training samples randomly, iteratively adding configurations for which the model has the highest absolute error results in a model which has a low error for predictions over the full range of different GPR standard deviations.
Figure 7: Correlation plots for GPR with Coulomb representations for Aspirin after 800 iterations of uncertainty sampling. The left plot in the first row shows the correlation plot with predicted GPR standard deviation vs. actual error for all configurations of the candidate pool of the initial GPR model that was trained with 200 random configurations. The right plot in the first row shows the uncertainty sampling results for 800 iterations. In the bottom row, correlation plots are shown for a GPR model trained with 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty estimation, as well as after iteratively adding 800 configurations with the highest absolute error and adding 800 additional training configurations randomly.
4 Discussion
The objective of this study was to evaluate different uncertainty estimators of GPR based MLIPs. To understand how uncertainty estimations correlate with the actual error, we plot both against each other and analyze if the errors are distributed according to the theoretical distributions, that result from the predictive distributions. For the GPR standard deviation, a clear correlation between an increasing uncertainty and increasing likelihood for larger errors is observed. For the two sets uncertainty, no correlation is observed, and the bootstrap uncertainty only shows a clear correlation in a few cases. However, even in cases where a correlation is observed, the actual error distributions often do not align with the error distributions that are predicted by the model. Therefore, the predictive distributions cannot quantify the error with valid confidence intervals. This also means it will not be possible to propagate uncertainty in a meaningful way. If the predictions of an MLIP are used to calculate further quantities, propagating uncertainty will not result in an accurate confidence intervals for the derived quantity. Still, if the uncertainty correlates with a higher likelihood for larger errors it provides a useful number, although a proper threshold might have to be found empirically. We found that specially predictions with higher uncertainties often will be highly overconfident, because the error is dominated by a high bias which is not captured quantitatively by the uncertainty. In contrast to our findings, Tran et al.38 found that the GPR standard deviation of their GPR model predicts underconfident confidence intervals for their exemplary dataset. The GAP model of Bartok et al.44 on the other hand predicts overconfident confidence intervals for their dataset of Ar dimers and trimers too, if the hyperparameters are optimized by maximizing the marginal likelihood, like we do. When optimizing the hyperparameters by optimizing the leave-one-out cross validation likelihood their GPR model predicts underconfident confidence intervals.
In ML theory, the error of a model is often separated into an irreducible part which results from the noise in the training data, as well as a bias and a variance part. We observe that predictions with a high uncertainty often have a high bias, which is not captured by the uncertainty in a quantitative way. Hüllermeier and Waegeman54 describe the GPR standard deviation as meaningful indicator for the total model uncertainty. However, if we have a model that has a bias in the form of systematical shifts of the predictive distributions in some areas of the input space it is not possible, that the standard deviation of a predictive distribution quantifies this shifts. Therefore, in general, a bias is associated with incorrect model assumptions, which makes it challenging to quantify the resulting error with the flawed model itself. For ensemble based uncertainties Heid et al.47 discussed that a bias can not be quantified.
When sampling a fixed configuration space with uncertainty sampling, and the GPR standard deviation is used as uncertainty estimator, the test MAE does not improve compared to random sampling in cases where the predictive distributions aligned well with the error distributions, while the maximum absolute error and the test variance sometimes improve compared to random sampling. For cases where the predictions with increasing uncertainty have an increasing bias, which are most of the cases, the MAE is getting worse compared to random sampling. This aligns well with the findings of Uteva et al.9, where uncertainty sampling with the GPR variance performs worse than uncertainty sampling with the two sets uncertainty and adding samples with the highest absolute error, as well as training with maximum Latin hypercube data. This behaviour might be explained by the fact that the GPR standard deviation finds a maximally distant training set, which results in a model that overemphasizes the border of the configuration space, that was originally sampled in the fixed dataset. The borders of the configuration space, which is represented in the fixed dataset might be sampled more sparse while other areas are sampled more densely. A model that overemphasizes the border might perform slightly worse on samples of other areas of the configuration space, which are more frequent. While this means that uncertainty sampling can result in an increased average model error inside the fixed configuration space, it is likely to be beneficial for extrapolation tasks. This observation was also made by Podryabinkin et al.7 when they trained an MLIP with the most uncertain configurations of a fixed dataset, according to the D-optimality criterium. Compared to random sampling, the model shows a slightly increased root mean squared error and a slightly decreased maximum error, when testing on a test dataset, which was previously selected randomly from the fixed dataset, compared to random sampling, while MD simulations with this model are much more stable. The reduced data efficiency, when sampling a fixed configuration space with uncertainty sampling, suggest that for on-the-fly AL, where uncertainty estimations guide the exploration of configuration space, data efficiency may also be compromised. Contrary to this, Vandermause et al.3 were showing an increased data efficiency with their on-the-fly AL scheme applied to 5-element high entropy alloy structures.
Note that our tests were performed for energy predictions without forces. Other studies have reported that uncertainty estimations based on force predictions can be beneficial, because as forces are the derivative with respect to the atom coordinates, they carry more local information. Besides that, in general a higher model accuracy also leads to better uncertainty estimations. For the SOAP representations, parameters, which lead to higher resolution of the atomistic environments, would be possible. Still, we already use an atomistic kernel without sparsification, which in general leads to a much more accurate model than the classical GAP model with sparsification33.
In cases where the error distributions align well with the predictive GPR distributions, we find that data efficiency does not improve. If error distributions align with the predictive GPR distributions, even the errors of predictions with a high uncertainty are distributed around zero and many configurations for which the model has a high uncertainty actually have a low error. When we add samples with the actual highest error, we observe a significantly improved data efficiency. Therefore, if we would have a more linear correlation, i.e. the error is directly proportional to the uncertainty, between uncertainty estimations and errors instead of a correlation between uncertainty and how broad error distributions are distributed around zero, we might archive a much better data efficiency. One approach to archive this is to average over atomistic uncertainties55. If we would have calculated the bin averages of the absolute error in our correlation plots instead of the error, for cases where the predictive GPR distributions aligned well with the error distributions we would see a linear correlation between uncertainty and absolute error. However, this linear correlation is only archived because we average over many samples, for a single sample the most likely absolute error is zero. For atomistic models, like GPR with SOAP, the energy prediction can be seen as a sum of atomistic energy predictions of atomistic submodels. If forces are considered the force prediction for every atom can be seen as an atomistic submodel too. Now averaging the uncertainty prediction of these atomistic submodels can lead to a linear correlation between uncertainty estimation and absolute error.
5 Conclusion
In this study, we evaluated how different uncertainty estimations for GPR based MLIPs relate to the actual error and if sampling a fixed configuration space by iteratively choosing configuration with the maximum uncertainty can improve the model accuracy compared to random sampling. We find that ensemble based uncertainties often do not provide meaningful information about the error, while for the GPR standard deviation a clear correlation between increasing uncertainty and increasing likelihood for higher errors is observed. Still also the GPR standard deviation for most datasets is not able to quantify the prediction error correctly, therefore no accurate confidence intervals for model predictions can be derived. While an increasing GPR standard deviation correlates with an increasing bias in the model predictions, this bias is not captured quantitatively by the uncertainty estimation and often dominates the error of predictions with high uncertainty, resulting in highly overconfident confidence intervals. When sampling a fixed configuration space with configurations for which a GPR model predicts maximum GPR standard deviation, training configurations are found which are maximally distant to each other. This may lead to a more general model, that performs better for extrapolation tasks. Still, such a model can overemphasize the borders of the configuration space of a fixed dataset and perform on average worse on the fixed dataset than a model based on random sampling.
Acknowledgment
The authors acknowledge support by the DFG through the DFG Priority Program SPP 2363 on “Utilization and Development of Machine Learning for Molecular Applications – Molecular Machine Learning” through the projects ZA 1175/4-1 and KL 1299/25-1. Some of the computations were carried out on the PLEIADES cluster at the University of Wuppertal, which was supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF). The authors would also like to acknowledge the support of the ‘Interdisciplinary Center for Machine Learning and Data Analytics (IZMD)’ at the University of Wuppertal.
S1 Hyperparameter optimization
To optimize the GPR hyperparameters we maximize the marginal log likelihood (MLL) using gpytorch45. We use the ADAM optimizer with a learning rate of $lr=0.1$ and 200 steps. Since prior assumption of the hyperparameters, i.e. the hyperparameter starting values in the optimization process, can significantly influence the outcome, we perform cross validation for all combinations of possible starting values shown in Table S2. In all cases we use 1000 random samples to optimize the hyperparameters, for which we perform 5-fold cross validation with 5 repetitions. In every fold we optimize the hyperparameters with the respective training samples and starting values. Then we choose the hyperparamter starting values which resulted in the lowest loss on average over all repetitions and folds and use them as starting values for maximizing the MLL with all 1000 random samples.
Hyperparameter Possible values
Lengthscales $l_{1},...,l_{D}$ 2, $10^{0.75}$, $10^{1.5}$
Outputscale $\sigma_{f}$ 1
Noise $\sigma^{2}_{n}$ $10^{-4}$, $10^{-6}$, $10^{-8}$
Table S1: Possible starting values for hyperparameters for the maximization of the MLL.
S2 Ensemble mean vs single model mean
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S1: Correlation plots of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset. On the one hand the predictions are the mean of the submodel predictions of the two sets or bootstrap ensemble, on the other hand the predictions are from a single GPR model, trained on the full training set, from which subsets are used to train the ensemble submodels.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S2: Correlation plots of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for aspirin from the rMD17 dataset. On the one hand the predictions are the mean of the submodel predictions of the two sets or bootstrap ensemble, on the other hand the predictions are from a single GPR model, trained on the full training set, from which subsets are used to train the ensemble submodels.
S3 Bin widths
Molecule Bin width in meV
rMD17 Benzene 1
rMD17 Aspirin 15
WS22 SMA 20
WS22 O-HBDI 100
Porphyrin 2
Table S2: Bin width of correlation plots for different molecules.
S4 Comparison of Uncertainty error correlation between several molecules
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S3: Correlation of the two sets uncertainty and error across different datasets for GPR with Coulomb and GPR with SOAP. The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical error distributions. Note that for porphyrin and GPR with Coulomb the axis ranges and and bin width are adapted in comparison to the correlation plot for the GPR standard deviation for better visibility.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S4: Correlation of the bootstrap uncertainty and error across different datasets for GPR with Coulomb and GPR with SOAP. The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical error distributions. Note that for porphyrin and GPR with Coulomb the axis ranges and and bin width are adapted in comparison to the correlation plot for the GPR standard deviation for better visibility.
S5 Comparison of uncertainty sampling between several molecules
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S5: Maximum absolute error of every uncertainty sampling iteration for GPR with Coulomb and SOAP representations for different datasets. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure S6: Test variance of every uncertainty sampling iteration for GPR with Coulomb and SOAP representations for different datasets. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error.

### v2

@@
 Abstract
-Machine learning interatomic potentials (MLIPs) have seen significant advances as efficient replacement of expensive quantum chemical calculations for predicting molecular energies and forces.
-Uncertainty estimations for these machine learning models are crucial to quantify the additional model error they introduce and to leverage this information to find most informative training samples in active learning strategies.
-MLIPs that are based on Gaussian process regression provide predictive distributions with a standard deviation as a possible uncertainty measure.
-An alternative approach are ensemble-based uncertainties.
-Although these uncertainty measures have been applied to active learning, it has rarely been studied how well or whether they correlate with the actual error, and it is not always clear whether active learning strategies actually outperform randomized sampling strategies.
-In our study, we consider GPR models with Coulomb and Smooth Overlap of Atomic Positions (SOAP) representations as inputs to predict potential energy surfaces and excitation energies of molecules.
-We evaluate, how the GPR variance and ensemble-based uncertainties relate to the actual error and whether model performance improves by selecting the most uncertain samples from a fixed configuration space for training.
-For the ensemble based uncertainty estimations, we find that they often do not provide any information about the error.
-For the predictive GPR standard deviation, we find that in most cases predictions with an increasing standard deviation also have an increasing systematical bias, which is not captured by the uncertainty.
-In these cases, selecting training samples with the highest uncertainty leads to a model with a worse test error compared to selecting samples randomly.
-We conclude that confidence intervals for GPR-based MLIPs, which are derived from the predictive standard deviation, can be highly overconfident, as they do not capture the bias in the model predictions.
-Selecting samples with high GPR standard deviation leads to a model that overemphasizes the borders of the configuration space represented in the fixed dataset.
+Uncertainty estimations for machine learning interatomic potentials (MLIPs) are crucial for quantifying model error and identifying informative training samples in active learning strategies.
+In this study, we evaluate uncertainty estimations of Gaussian process regression (GPR)-based MLIPs, including the predictive GPR standard deviation and ensemble-based uncertainties.
+We do this in terms of calibration and in terms of impact on model performance in an active learning scheme.
+We consider GPR models with Coulomb and Smooth Overlap of Atomic Positions (SOAP) representations as inputs to predict potential energy surfaces and excitation energies of molecules.
+Regarding calibration, we find that ensemble-based uncertainty estimations show already poor global calibration (e.g., averaged over the whole test set).
+In contrast, the GPR standard deviation shows good global calibration, but when grouping predictions by their uncertainty, we observe a systematical bias for predictions with high uncertainty.
+Although an increasing uncertainty correlates with an increasing bias, the bias is not captured quantitatively by the uncertainty.
+Therefore, the GPR standard deviation can be useful to identify predictions with a high bias and error but, without further knowledge, should not be interpreted as a quantitative measure for a potential error range.
+Selecting the samples with the highest GPR standard deviation from a fixed configuration space leads to a model that overemphasizes the borders of the configuration space represented in the fixed dataset.
 This may result in worse performance in more densely sampled areas but better generalization for extrapolation tasks.
@@
 These samples then serve as training data for a machine learning model, which can predict new samples at substantially lower costs.
-How well samples from different parts of the configuration space are predicted by an MLIP depends highly on its training data.
-By configuration space, we refer to the set of all possible spatial arrangements of atoms in an atomic system.
-Samples from areas of the configuration space that differ too much from those covered by the training data might not be predicted well.
-Because of this variability in prediction accuracy, it becomes desirable to provide an uncertainty estimation for every model prediction, which ideally gives a good estimation for the prediction error.
-A good error estimation could help to understand the range of a potential error of a prediction, and could be utilized in active learning (AL) strategies1; 2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15; 16; 17.
+How well samples from different areas of the configuration space are predicted by an MLIP depends highly on its training data.
+Samples from areas of the configuration space (e.g., spatial arrangements of atoms in an atomic system) that differ too much from those covered by the training data might not be predicted well.
+Because of this variability in prediction accuracy, it becomes desirable to provide an uncertainty estimation for every model prediction.
+A good uncertainty estimation could help to understand the range of a potential error of a prediction, and be utilized in active learning (AL) strategies[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
 With AL, algorithms are referred to that aim to choose or create new training samples, which are maximally informative to the model.
-In the context of MLIPs, two important goals of AL exist.
-The first goal is to increase the data efficiency of the model training.
-Data efficiency refers to the reduction in the number of training samples needed to achieve comparable model accuracy within a fixed configuration space, compared to using randomly selected samples.
-A second goal is to sample the configuration space, without being fixed, in all areas which are relevant for the respective applications.
 Many different ML approaches to design MLIPs exist.
-One class of MLIPs are message passing neural networks 18; 19; 20; 21 (MPNN), which contain features that are trained via message passing steps.
-In other approaches, instead of trained features, fixed representations are employed that either globally represent the whole system of atoms 22; 23; 24 or locally represent atomistic environments 25; 26; 27; 28.
-These representations provide important properties such as invariance against translation, rotation, and atom permutation29.
-Besides linear models 26; 30 and neural networks28; 31; 32, kernel-based models are commonly chosen as ML model 33; 22; 24.
-Within kernel-based models, many are variations of Gaussian process regression34 (GPR).
-GPR predicts the output for a new input by generating a Gaussian predictive distribution, where the mean represents the (deterministic) prediction and the variance provides a measure of uncertainty for the prediction.
-The GPR variance is commonly used as an uncertainty measure in AL schemes for GPR-based MLIPs9; 16; 3; 1.
-Besides the GPR variance, other uncertainty estimations are conceivable for GPR.
-A common approach in ML is to construct ensembles of slightly different models and use the variance of their predictions as an uncertainty measure.
-Model ensembles can be based on different initializations of weights 35 or they can be generated by training on different subsets of the training data36.
-For neural networks, using different weight initializations is reasonable because these models are parametric and sensitive to their initial conditions.
-In contrast, for GPR, it is more reasonable to use different subsets of the training data because GPR is a non-parametric method that relies directly on the training data to make predictions.
-Uteva et al.9 trained two models on half of the initial training set each.
-Similarly, a bootstrap aggregation36; 37 ensemble, which is commonly used in ML, would be possible for GPR.
-Here multiple models are trained on different subsets of the training data, which are generated by randomly sampling indices from the original dataset.
-Our focus will be on GPR-based MLIPs with GPR variance and ensemble approaches as uncertainty measure or predictive distributions.
-GPR models, as well as GPR ensembles, provide predictive distributions with a mean and a standard deviation.
-The mean of this distribution may serve as point prediction, which is the model’s actual predicted value.
-The standard deviation is a measure of variability or uncertainty in the prediction.
-An important question is, to which extend such uncertainty measures are correlated to the prediction error of a model, thus whether they would serve as error estimator.
-This requires a careful evaluation.
-Evaluating uncertainty estimations of MLIPs is a challenging problem with a large variety of approached being reported in the literature.
-Tran et al.38 distinguished between three aspects of uncertainty estimation: calibration, sharpness and dispersion.
-Calibration describes if the uncertainty estimation is correct, in the sense that the true error values follow the predicted distribution.
-Sharpness characterizes the overall confidence of a model’s predictions, and generally, a model with more confident predictions is preferred.
-Dispersion, on the other hand, refers to the spread of the uncertainty estimates.
-A model with high dispersion shows a wide variability in confidence across different predictions, allowing it to differentiate between samples predicted with high accuracy and those potentially having high errors.
-Tran et al.38 were analyzing these aspects with a variety of metrics.
-In other studies, often the uncertainty estimations are plotted against the actual or absolute errors3; 39; 40; 41; 10, which, in principal, gives an intuitively accessible picture off all three aspects.
-Kellner et al.
-39 constructed an example with synthetic data to show how an ideal uncertainty versus error double-logarithmic plot would look like and illustrated various cases of incorrect uncertainty estimations.
-They then analyzed, if uncertainty versus error plots for non-synthetic data show the correct shape.
-Busk et al.
-40 sorted all predictions in equally sized bins and calculated metrics like the z-score variance and the RMSE for all samples of one bin.
-In this work, we use binning to analyze the error distributions of predictions with certain uncertainties.
-In the literature, uncertainty estimations of MLIPs are applied in a large variety of different active learning schemes.
-Many of these create new configurations without explicitly restricting the configuration space.
-A particularly important AL scheme is on-the-fly AL.
-Here, an uncertainty is calculated at every point during an MD run with the MLIP to recognize if the MLIP runs into configuration areas which are poorly sampled in the training data and for which the MD run might become instable1; 2; 3; 4; 5; 6; 7.
-Other schemes are optimizing the uncertainty estimation multiplied by a weight function to create new training configurations15; 16; 17.
-However, when using uncertainty estimation to guide the exploration of a flexible configuration space, it becomes challenging to assess the actual impact of active learning and the chosen uncertainty metric in comparison to generating samples randomly.
-Therefore, we restrict ourselves to a simple, very common AL approach from standard ML literature, which samples a fixed configuration space.
-This approach is often referred to as uncertainty sampling 42; 9.
-Uncertainty sampling can be applied in setups, where a large amount of unlabeled data exists and generating the respective labels is expensive.
-For a small random subset of the data, labels are generated.
-Then this data is used to train an initial model.
-The uncertainty estimation of this model is used to iteratively select samples from the remaining pool of unlabeled samples for which the labels are generated and which are then added as additional training samples.
-In the general ML literature, uncertainty sampling is often applied to classification problems, but it is also used for regression problems 43.
-It was shown that uncertainty sampling can increase the data efficiency by a multiple, although the effect strongly depends on the dataset and the uncertainty estimation that is used.
-In this study, uncertainty sampling is applied to analyze the benefit of the discussed different uncertainty estimation methods on AL.
-In literature, uncertainty estimations of GPR based MLIPs are often evaluated by directly examining the performance when applied to AL without explicitly analyzing how the uncertainty estimations correlate with the actual error9; 1; 16.
-These studies give a mixed picture on the effect of AL on training GPR-based MLIPs and it is often not clear, whether AL strategies provide an actual improvement over random sampling.
-The correlation between uncertainty estimations of GPR-based MLIPs and the actual error has rarely been studied, explicitly.
-Vandermause et al.3 plotted their Gaussian process uncertainty against the true absolute error and reported that most true absolute errors fall within twice the Gaussian process uncertainty.
-They further showed how their active learning method, based on the Gaussian process uncertainty, provides actual improvement over random sampling.
-Bartok et al.
-44 plotted uncertainty estimations against the true absolute error for a dataset of Ar dimers and trimers, finding that maximizing the leave-one-out cross-validation likelihood for hyperparameter optimization can be advantageous for uncertainty estimation, compared to maximizing the marginal likelihood or manual hyperparameter tuning.
-Tran et al.
-38 evaluated different aspects of uncertainty estimations for an exemplary dataset using, among others, a GPR-based MLIP.
-Various GPR-based MLIPs exist, which differ in how their kernel is constructed and which representations are used.
-Therefore, also the GPR variance differs and it is not clear how the GPR variances of different models compare as uncertainty estimations.
-Furthermore, even if an uncertainty estimation shows a meaningful correlation with the errors, translating this relationship into enhanced data efficiency for AL runs remains uncertain.
-The objective of this study is hence to evaluate uncertainty measures for GPR-based MLIPs.
-We consider a GPR model with global Coulomb representations22 as inputs as well as an atomistic GPR model with SOAP25 representations.
-We test the prediction of ground state energies from molecules of well-established benchmarks, allowing for easy comparison and categorization of our results.
-We further predict excited state energies for a trajectory of porphyrin in gas phase as a large and complex molecule with a potentially complex configuration space.
-To evaluate the uncertainty estimations, we test how the GPR variance, as well as ensemble uncertainties, of different GPR-based MLIPs, correlate with the actual error for unseen samples.
-This is done by analyzing the distributions of errors of predictions with different uncertainties.
-It is further investigated how different uncertainty measures perform, when applied to AL.
-Uncertainty sampling is chosen as AL algorithm for the reasons discussed above.
-To understand, how much information could be maximally extracted from the candidate pool, when using uncertainty sampling, we always also provide a reference AL experiment.
-In that reference AL experiment, exact information on the error is used, in contrast to using an uncertainty measure.
-This work is structured as follows.
-In Section 2, Gaussian Process Regression (GPR) is introduced, and the use of Coulomb and SOAP representations to construct GPR-based MLIPs is described.
-The different uncertainty estimations that are applied are described and we give a detailed description of how we evaluate them.
-Further details are given on how uncertainty sampling was applied and which data is used.
-In Section 3 we analyze the correlation of uncertainty estimations and errors for the different datasets.
-It is further investigated how the observed correlations transfer when applied to uncertainty sampling.
-An overall discussion of the results is given in Section 4.
-Conclusions with an outlook are given in Section 5.
+One class of MLIPs are message passing neural networks [18, 19, 20, 21] (MPNN), which contain features that are trained via message passing steps.
+In other approaches, instead of trained features, fixed representations are employed that either globally represent the whole system of atoms [22, 23, 24] or locally represent atomistic environments [25, 26, 27, 28].
+These representations provide important properties such as invariance against translation, rotation, and atom permutation[29].
+Besides linear models [26, 30] and neural networks[28, 31, 32], kernel-based models are commonly chosen as ML model [33, 22, 24].
+Within kernel-based models, many are variations of Gaussian process regression[34] (GPR).
+GPR predicts the output for a new input by generating a Gaussian predictive distribution, where the mean represents the (deterministic) prediction, and the standard deviation provides a measure of uncertainty for the prediction.
+The GPR standard deviation is commonly used as an uncertainty measure in AL schemes for GPR-based MLIPs[9, 16, 3, 1].
+Besides the GPR standard deviation, other uncertainty estimations are conceivable for GPR.
+A common approach in ML is to construct ensembles of slightly different models and use the standard deviation of their predictions as an uncertainty measure.
+Since GPR is a non-parametric method that relies directly on the training data for predictions, it is reasonable to construct ensembles where the models differ based on subsets of the training data.
+Uteva et al.[9] trained two models on half of the initial training set each.
+Similarly, a bootstrap aggregation[35, 36] ensemble, which is commonly used in ML, can be applied to GPR.
+Both the GPR standard deviation and ensemble-based uncertainties are standard deviations of predictive distributions, thus provide uncertainty estimations.
+To evaluate these uncertainty estimations, the most important aspect is calibration, which describes how well the distribution of empirical errors aligns with the predictive distributions.
+A common tool to evaluate calibration are calibration curves[37], which test whether the empirical errors are covered by the confidence intervals derived from the predictive distributions.
+However, it has been pointed out that this approach does not sufficiently validate good calibration, because it averages over the entire test dataset [38, 39].
+Separating predictions into bins by their uncertainty can give a more localized evaluation of uncertainty estimations, for example, using reliability diagrams[39], or by calculating the variance of z-scores[38].
+In the literature evaluating uncertainty estimations of general MLIPs, local analyses of calibration are often lacking or limited to qualitative tests, such as plotting errors against uncertainties [40, 41, 10, 42].
+In contrast, Busk et al.
+[43] sorted all predictions into bins to generate reliability diagrams and calculate local z-score variances.
+In this work, we use calibration curves as well as extended reliability diagrams to evaluate calibration.
+While classical reliability diagrams only show the standard deviation of empirical errors within each bin, extended reliability diagrams additionally show the mean of empirical errors.
+This accounts for the fact that, when only the standard deviation of errors is shown, a systematical shift of the error distribution is not visible.
+Beside explicitly analyzing calibration, uncertainty estimations can also be evaluated by their performance when utilized in AL schemes.
+While a wide variety of AL schemes exist in the MLIP literature, in this work we restrict ourselves to a simple and widely used AL approach from standard machine learning literature, known as uncertainty sampling [44, 9].
+Uncertainty sampling samples a fixed configuration space, represented in a fixed dataset.
+An initial model is trained on a small random subset of the dataset, and the uncertainty estimation of this model is then used to iteratively select samples from the remaining pool.
+In the general ML literature, uncertainty sampling is often applied to classification problems, but it is also used for regression problems [45].
+It was shown that uncertainty sampling can improve data efficiency, which refers to the reduction in the number of training samples needed to achieve comparable model accuracy compared to using randomly selected samples, by a multiple.
+However, the extent of this improvement strongly depends on the dataset and the uncertainty measure used [45].
+In this study, uncertainty sampling is applied to analyze the benefit of the discussed different uncertainty measures on AL.
+When it comes to GPR-based MLIPs, the calibration of uncertainty estimations has rarely been evaluated explicitly.
+Aside from qualitative visualizations of calibration, such as plotting errors against uncertainties[3, 46], Tran et al.
+[40] evaluated different aspects of uncertainty estimations for an exemplary dataset using, among others, a GPR-based MLIP.
+More often, uncertainty estimations of GPR-based MLIPs are evaluated by directly examining the performance when applied to AL, without explicitly analyzing calibration[3, 9, 1, 16].
+However, these studies give a mixed picture on the effect of AL on training GPR-based MLIPs, and it is not always clear whether AL strategies provide an actual improvement over random sampling.
+The objective of this study is to provide a comprehensive analysis of uncertainty estimations for GPR-based MLIPs.
+To achieve this, we provide an explicit evaluation of calibration and investigate the performance of different uncertainty measures applied to active learning (AL).
+We analyze global calibration using calibration curves and local calibration of groups of predictions, separated by their uncertainty, with extended reliability diagrams.
+Uncertainty sampling is used as a simple AL algorithm.
+We consider a GPR model with global Coulomb representations[22] as inputs, as well as an atomistic GPR model with SOAP[25] representations.
+Our tests include predicting ground-state energies for molecules from well-established benchmarks.
+Additionally, we predict excited-state energies for a trajectory of porphyrin in the gas phase, chosen as a large and complex molecule with a potentially complex configuration space.
 2 Methods
 2.1 Gaussian process regression
-In GPR34, a prior probability distribution over functions is defined using a Gaussian process.
+In GPR[34], a prior probability distribution over functions is defined using a Gaussian process.
 We want to find a posterior predictive distribution from training inputs $\mathbf{x}_{i}$ and corresponding output measurements $y_{i}$, with which we can predict outputs for unknown inputs x.
@@
 The GPR framework provides a predictive distribution for the output corresponding to any input x.
-The point prediction for the output of an input $\mathbf{x}$ is the mean of this predictive distribution, given as
+The actually predicted value for the output of an input $\mathbf{x}$ is the mean of this predictive distribution, given as
 $m\left(\mathbf{x}\right)=\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{y}\,.$ (2)
@@
 which is one of the most common choices.
-We used a different length-scale for every feature, known as automatic relevance determination34, to increase the accuracy and flexibility of the model.
+We used a different length-scale for every feature, known as automatic relevance determination[34], to increase the accuracy and flexibility of the model.
 Here $\mathbf{M}$ is a diagonal matrix $\text{diag}\left(\frac{1}{l_{1}^{2}},...,\frac{1}{l_{D}^{2}}\right)$, with $D$ the number of features, and $\sigma_{f}^{2}$ is the output-scale.
-We optimized the hyperparameters $\sigma_{f}^{2}$ and $\sigma_{n}^{2}$, as well as length-scales inside the covariance function via maximum marginal log-likelihood, details can be found in Section S2.
-For all our tests, we used the GPR implementation of GPyTorch 45.
+We optimized the hyperparameters $\sigma_{f}^{2}$ and $\sigma_{n}^{2}$, as well as length-scales inside the covariance function via maximum marginal log-likelihood, details can be found in the section hyperparameter optimization of the supplementary material.
+For all our tests, we used the GPR implementation of GPyTorch [47].
 In the case of a global representation as inputs $\mathbf{x}$, the covariance function of Eq.
 3 is directly applied to these inputs.
-For our study, we incorporate Coulomb matrices22 as a global representation.
-For Coulomb matrices, to ensure invariance against atom indexing, often a sorting is carried out, which results in discontinuities, which may cause challenges in the construction of ML models 29.
-In our tests, we ignored invariance against atom permutation and omitted sorting because, for all tasks, a system with fixed atoms given in a fixed order was given.
-As a local representation, we used Smooth overlap of atomic positions (SOAP) 25.
-It is commonly used and GPR models with SOAP representations as inputs have shown high accuracy 29.
+For our study, we incorporated Coulomb matrices[22] as a global representation.
+For Coulomb matrices, to ensure invariance against atom indexing, often a sorting is carried out, which results in discontinuities, which may cause challenges in the construction of ML models [29].
+In our tests, we ignored invariance against atom permutation and omitted sorting because, for all tasks, a system with fixed atoms in a fixed order was given.
+As a local representation, we used Smooth overlap of atomic positions (SOAP) [25].
+It is commonly used and GPR models with SOAP representations as inputs have shown high accuracy [29].
 For atomistic representations, an input x will consist of representations $\textbf{x}_{j}$ for every atomistic environment of the system.
@@
 In fact, the predictive distribution of a GPR model is anyway Gaussian.
-Moreover, the predictive distributions of GPR ensembles are often assumed to be Gaussian too40; 47, which is, according to the central limit theorem, at least true for a large number of ensemble models47.
+Moreover, the predictive distributions of GPR ensembles are often assumed to be Gaussian too[43, 49], which is, according to the central limit theorem, at least true for a large number of ensemble models[49].
 For the case of the standard GPR model, the predictive distribution is, using notation from eq. (5), given by a pair
 $\mathcal{D}_{G}(m_{G}(\mathbf{x}),u_{G}(\mathbf{x})^{2})\,,$ (6)
-where $m_{g}(\mathbf{x})$ is the GPR mean from eq. 2 and we have
-$u_{g}\left(\textbf{x}\right)=\sigma\left(\textbf{x}\right)$ (7)
-for all $\mathbf{x}$, with the GPR variance
-$\sigma^{2}\left(\mathbf{x}\right)=k\left(\mathbf{x},\mathbf{x}\right)-\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{K}\left(\mathbf{X},\mathbf{x}\right)\,.$ (8)
+where $m_{G}(\mathbf{x})$ is the GPR mean from eq. 2 and we have the GPR standard deviation
+$u_{G}\left(\textbf{x}\right)=\sqrt{k\left(\mathbf{x},\mathbf{x}\right)-\mathbf{K}\left(\mathbf{x},\mathbf{X}\right)\bigl{(}\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{K}\left(\mathbf{X},\mathbf{x}\right)}\,.$ (7)
 Note that the GPR variance is only calculated from the kernel function $k(\cdot,\cdot)$, the inputs $\mathbf{X}$ of the training set, i.e.
@@
 In the variance, see eq.
-8, GPR already accounts for the noise $\sigma_{n}^{2}$ in the observations, therefore, we do not add this noise to our uncertainty estimation, like it was done by Vandermause et al.3.
-The two set approach, proposed by Uteva et al.9, was motivated by the principles underlying the jackknife and bootstrap methods36. We use for its predictive distribution the notation
+(7), GPR already accounts for the noise $\sigma_{n}^{2}$ in the observations, therefore, we do not add this noise to our uncertainty estimation, like it was done by Vandermause et al.[3].
+The two set approach, proposed by Uteva et al.[9], was motivated by the principles underlying the jackknife and bootstrap methods[35]. We use for its predictive distribution the notation
 $\mathcal{D}_{2}(m_{2}(\mathbf{x}),u_{2}(\mathbf{x})^{2})\,.$ (8)
@@
 The square of this particular uncertainty measure can be interpreted as a very rough estimate of the model’s variance with respect to changes in the training set, including the energies.
-In other words, it very roughly approximates the variance in the bias-variance decomposition of ML models, which is common in ML literature48.
-An extension of this is the bootstrap aggregation ensemble36, which leads to the predictive distribution
+In other words, it very roughly approximates the variance in the bias-variance decomposition of ML models, which is common in ML literature[50].
+An extension of this is the bootstrap aggregation ensemble[35], which leads to the predictive distribution
 $\mathcal{D}_{b}=(m_{b}(\mathbf{x}),u_{b}(\mathbf{x})^{2})\,.$ (11)
@@
 The mean is then
-${m}_{b}\left(\textbf{x}\right)=\sum_{l=1}^{N}m^{\left(l\right)}\left(\textbf{x}\right)\,,$ (13)
+${m}_{b}\left(\textbf{x}\right)=\frac{1}{N}\sum_{l=1}^{N}m^{\left(l\right)}\left(\textbf{x}\right)\,,$ (12)
 and the uncertainty is calculated by the standard deviation of their different predictions $m^{(l)}\left(\textbf{x}\right)$,
@@
 As a slight change towards the classical definition of the two-sets and bootstrap approach, we do not use the ensemble mean as mean for their predictive distributions, since we found whether we use the mean of the ”full” GPR model or the ensemble mean does not make a difference for the essential results on the correlation between error and uncertainty estimation.
-A comparison can be found in Section S2.
+A comparison can be found in Fig.
+S1 and S2 of the supplementary material.
 2.3 Evaluating uncertainty estimations
-The main objective of this study is to analyze the relationship of uncertainty estimations, such as GPR variance, and the actual prediction error.
-In particular, it is of interest to understand whether uncertainty estimations of ML models can serve as error estimator.
-A straight forward way to analyze this is to calculate uncertainty estimations and errors of a set of $n_{\text{test}}$ unseen test configurations and plot the uncertainty estimations against the errors.
-To better understand, what we could potentially expect from uncertainty estimations in an ideal setting, let us assume we would have an approach that would construct an optimal predictive distribution
+The most important aspect to analyze uncertainty estimations, such as GPR standard deviation, is their relation to the actual prediction error.
+To better understand, what we could potentially expect from uncertainty estimations in an ideal setting, let us assume an approach that constructs an optimal predictive distribution
 $\mathcal{D}_{opt}(m_{opt}(\mathbf{x}),u_{opt}(\mathbf{x})^{2})\,,$
-where the true values $y(\mathbf{x})$ related to predictions $m_{opt}(\mathbf{x})$ are actually following the distribution
+where the true values $y(\mathbf{x})$ corresponding to predictions $m_{opt}(\mathbf{x})$ actually follow the distribution
 $y(\mathbf{x})\sim\mathcal{N}(m_{opt}(\mathbf{x}),u_{opt}(\mathbf{x})^{2})\,.$
-It is then quite immediate that, under this thought experiment, the prediction error $\epsilon(\mathbf{x})$ of the model for all test samples would behave like
+It is then quite immediate that the prediction error $\epsilon(\mathbf{x})$ of the model for all test samples would behave like
 $\epsilon(\mathbf{x})=y(\mathbf{x})-m_{opt}(\mathbf{x})\sim\mathcal{N}(0,u_{opt}(\mathbf{x})^{2})\,.$
 In other words, the error of the model would be normally distributed with a mean of 00 and a standard deviation of $u_{opt}(\mathbf{x})$.
-In principle, this what we assume, if we consider the standard deviation of the predictive distribution to match, or to be perfectly correlated, to the error of the model.
 Consequently, our subsequent evaluation of the uncertainty estimations aims at testing for a given uncertainty measure $u(\mathbf{x})$, how closely the prediction errors $\epsilon$ would fulfill
 $\epsilon\sim\mathcal{N}(0,u(\mathbf{x})^{2}))\,.$ (14)
-If the observed error distribution closely aligns with eq.
-(15), valid confidence intervals, which provide a range that is expected to contain the true error with a specified probability, can be derived from the uncertainties of predictions.
-To validate if the errors of predictions with the same predictive variance are distributed according to eq.
-(15), we can separate all predictions by their uncertainty values into $n_{\text{bins}}$ equidistant bins with intervals
-$I_{\beta}=\left[\left(\beta-1\right)\Delta u,\,\beta\Delta u\right),\qquad\text{with}\qquad\beta=1,...,n_{\text{bins}},$ (16)
-where $\Delta u$ is the bin width. We then assume that all predictions inside one bin approximately have the center of that bin
+The extent to which the empirical distribution of prediction errors aligns with this ideal distribution is referred to as calibration.
+A common method for evaluating calibration is the use of calibration curves introduced by Kuleshov et al.[37].
+For every prediction within a test set, the uncertainty estimation is used to calculate a confidence interval, providing a range expected to contain the true value with a specific probability $\alpha_{\text{predicted}}$.
+By iterating over different $\alpha_{\text{predicted}}$ values from 0 to 1, the proportion of errors $\alpha_{\text{observed}}$ falling within the corresponding confidence intervals is calculated and plotted against $\alpha_{\text{predicted}}$.
+We use the implementation provided in the uncertainty toolbox[51] to generate calibration curves.
+A limitation of calibration curves is that they evaluate only global calibration over the entire test set, without accounting for local calibration, e.g., differences in calibration for specific groups of predictions.
+To address this, we employ a method similar to the reliability diagrams proposed by Levi et al.[39].
+Predictions are grouped based on their uncertainty values into $n_{\text{bins}}$ bins with intervals
+$I_{\beta}=\left[\left(\beta-1\right)\Delta u,\,\beta\Delta u\right),\qquad\text{with}\qquad\beta=1,...,n_{\text{bins}}.$ (15)
+Different to Levi et al., we use equidistant bins, with a fixed bin size $\Delta u$, and assume that all predictions inside one bin approximately have the center of that bin
 $u_{\beta}(\mathbf{x})=\left(\beta-\frac{1}{2}\right)\Delta u$ (16)
 as same fixed uncertainty.
-To evaluate whether the true error distributions align with the model’s prediction, one can calculate the bin mean and bin standard deviation of the errors $\epsilon_{\beta}$ within each bin.
-The mean of $\epsilon_{\beta}$ should be zero and the standard deviation should be $u_{\beta}\left(\textbf{x}\right)$, according to eq.
+To evaluate whether the true error distributions align with the model’s predictions, we calculate the bin standard deviation of the errors $\epsilon_{\beta}$ within each bin and compare it with $u_{\beta}\left(\textbf{x}\right)$, which is the predicted standard deviation according to eq.
 (14).
-If the bin mean and bin standard deviation align with the model prediction, and the errors are also distributed according to a Gaussian, as it is predicted, also the confidence intervals, which are derived from the predictive distributions would be correct.
-We will refer to plots in which we plot the uncertainty estimation against the errors and mark the bin means and standard deviations of errors as correlation plots.
-To illustrate a model, which perfectly aligns with eq.
+This evaluates whether the errors are as broad or narrow as predicted, but not if they are systematically shifted.
+In contrast to Levi et al., we therefore also calculate the bin mean of the errors $\epsilon_{\beta}$ within each bin.
+If the error mean differs from zero, this indicates a systematical shift of errors and suggests that the predictive distribution is shifted with a constant bias compared to the true distribution of targets.
+We refer to plots where uncertainty estimations are plotted against the errors, with bin means and standard deviations of errors marked, as extended reliability diagrams.
+To illustrate a model perfectly aligned with eq.
 (14), Fig.
-1 shows a correlation plot for data, where the targets were drawn from the predictive distribution of the GPR model.
-The distribution of errors can differ from this case, and therefore from Eq.
-15, in two ways: the calculated mean of the errors is not zero and the calculated standard deviation of the errors is not $u_{\beta}\left(\textbf{x}\right)$.
-If the error mean differs from zero this indicates that the predictive distribution is shifted with a constant bias compared to the true distribution of targets.
-If the error standard deviation is smaller or larger than $u_{\beta}\left(\textbf{x}\right)$ it indicates, that the predictive distribution is either to narrow or broad, respectively.
-Figure 1: Correlation plot for sinus data with synthetic GPR test targets.
-To generate the data a GPR model was trained with noisy data points from a sinus function.
+1 shows an extended reliability diagrams for data where the targets were drawn from the predictive distribution of the GPR model.
+Figure 1: Extended reliability diagram for sinus data with synthetic GPR test targets.
+To generate the data, a GPR model was trained with noisy data points from a sinus function.
 Test targets where drawn from the predictive distributions of the GPR model.
 Thereby data is generated for which the GPR model perfectly captures the underlying data-generating process.
-For all test samples the predicted GPR standard deviation is plotted against the actual error of the prediction.
+For all test samples, the predicted GPR standard deviation is plotted against the actual error of the prediction.
 All predictions are separated by their standard deviation into equidistant bins.
@@
 2.4 Uncertainty sampling
-With correlation plots, we can get a picture of different aspects that determine the quality of uncertainty estimations.
-As an additional study, we utilize the uncertainty estimations in an AL scheme, namely uncertainty sampling 42 to see, how effective is these estimations are in practical use.
-In uncertainty sampling, a small random subset of data is selected, and labels are generated for this subset.
+Beside explicitly evaluating calibration, we evaluate uncertainty estimations when utilized in an AL scheme, namely uncertainty sampling [44], to see, how effective these estimations are in practical use.
+In uncertainty sampling, a small random subset of a fixed dataset is selected, and labels are generated for this subset.
 In our case, these labels are obtained via quantum chemical calculations of potential energies.
@@
 See Algorithm 1 for pseudocode.
-To analyze, how much uncertainty sampling with different uncertainty estimators increases data efficiency, we kept a test set separate from the data pool.
+To analyze, how much uncertainty sampling with different uncertainty measures increases data efficiency, we kept a test set separate from the data pool.
 In every iteration, we calculated the mean absolute error (MAE) for the test samples.
@@
 If, instead of adding one sample per iteration, batches of samples are added, there is a risk of adding multiple very similar samples, for all of whom the model has a high uncertainty.
-Although methods exist that balance informativeness, diversity, and representativeness11; 12, we do not use batch active learning approaches.
+Although methods exist that balance informativeness, diversity, and representativeness[11, 12], we do not use batch active learning approaches.
 Our goal to isolate the influence of the uncertainty estimation and get an accurate picture of how much additional information is extracted with different uncertainty estimations.
 Algorithm 1 Uncertainty Sampling
-0:  Unlabeled data pool $U$, initial sample size $n_{\text{init}}$, number of iterations $n_{\text{iter}}$
-1:  Randomly select $n_{\text{init}}$ samples from $U$ and obtain their labels to form initial labeled dataset $L$
-2:  $U$ = $U$$L$
+0:  Unlabeled data pool $U_{\text{pool}}$, initial sample size $n_{\text{init}}$, number of iterations $n_{\text{iter}}$
+1:  Randomly select $n_{\text{init}}$ samples from $U_{\text{pool}}$ to get $U_{\text{init}}$ and obtain their labels to form initial labeled dataset $L$
+2:  $U_{\text{pool}}$ = $U_{\text{pool}}$$U_{\text{init}}$
 3:  Train initial model $m$ using labeled data $L$
 4:  for $i=1$ to $n_{\text{iter}}$ do
-5:     Compute uncertainty for each sample in $U$ using model $m$
-6:     Select sample $x^{*}$ from $U$ with highest uncertainty
+5:     Compute uncertainty for each sample in $U_{\text{pool}}$ using model $m$
+6:     Select sample $x^{*}$ from $U_{\text{pool}}$ with highest uncertainty
 7:     Obtain label $y^{*}$ for selected sample $x^{*}$
-8:     $L=L\,\cup\,(x^{*},y^{*})$
-9:     $U=U$$x^{*}$
+8:     $L=L\,\cup\,\{(x^{*},y^{*})\}$
+9:     $U_{\text{pool}}=U_{\text{pool}}$$\{x^{*}\}$
 10:     Retrain model $m$ using updated labeled dataset $L$
@@
 2.5 Data
-MD1749 is a benchmark dataset of long molecular dynamic trajectories, calculated with density functial theory (DFT), for 10 small molecules.
-The revised MD17 (rMD17) dataset50 recalculated the energies as the original values were found to be noisy.
-Additionally the total number of snapshots from the separate trajectory was reduced to a selection of less correlated snapshots.
+MD17[52] is a benchmark dataset of long molecular dynamic trajectories, calculated with density functional theory (DFT), for 10 small molecules.
+Christensen et al.[53] recalculated the energies as the original values were found to be noisy.
+Additionally, the total number of snapshots from the separate trajectory was reduced to a selection of less correlated snapshots.
+These changes result in the revised MD17 (rMD17) dataset.
 We selected the rMD17 dataset for our tests because it is a well-established benchmark, allowing for easy comparison and categorization of our results.
 We performed tests for potential energies of Benzene, as a highly symmetric molecule and Aspirin, as a more complex molecule.
-The WS22 dataset51 consists of distributions of configurations obtained via Wigner sampling for 10 different organic molecules, which aims to cover an extensive configuration space.
-This is important for our tests to understand how uncertainty estimations can distinguish between areas of configuration spaces that are predicted accurately and those that are predicted poorly across a wide range of configurations.
-We performed tests for potential energies of SMA and O-HBDI.
+The WS22 dataset[54] consists of distributions of configurations obtained via Wigner sampling for 10 different organic molecules, which aims to cover an extensive configuration space.
+This is important for our tests to understand how uncertainty estimations can distinguish between areas of configuration spaces where predictions are accurate, and regions where predictions are poor across a wide range of configurations.
+We performed tests for potential energies of 2-(methyliminomethyl)phenol (SMA) and 4-(2-hydroxybenzylidene)-1,2-dimethyl-1H-imidazol-5(4H)-one (O-HBDI).
 Additionally, we considered porphyrin as a large and complex molecule with a potentially complex configuration space.
 The dataset consists of 100,000 snapshots of a quantum molecular dynamic trajectory calculated with Density Functional Tight Binding (DFTB).
-To calculate the DFTB trajectory we took the initial structure from ChemSpider52 (ID: 4086).
-After DFT (B3LYP/6-31G) geometry optimization, 150,000 fs ground state dynamics at 300 K were performed with the DFTB+ package53, and 100,000 frames were extracted from the end of the trajectory with a stride of 1 fs.
-We performed our tests for the corresponding excitation energies which we calculated with time-dependent long-range corrected DFTB (TD-LC-DFTB).
+To calculate the DFTB trajectory we took the initial structure from ChemSpider[55] (ID: 4086).
+After DFT (B3LYP/6-31G) geometry optimization, 150,000 fs ground state dynamics at 300 K were performed with the DFTB+ package[56], and 100,000 frames were extracted from the end of the trajectory with a stride of 1 fs.
+We performed our tests for the corresponding excitation energies, which we calculated with time-dependent long-range corrected DFTB (TD-LC-DFTB).
+Figure 2 gives the chemical structures of all molecules used in this study.
 Figure 2: Chemical structures of the molecules used in the tests.
@@
 For every dataset, we use the 1000 training samples to optimize the hyperparameters.
-To generate the correlation plots, we calculate the uncertainty estimations and prediction errors of the respective GPR model for the candidate pool samples and plot them against each other.
-We separated the samples into equidistant bins by their uncertainty estimations and calculated the bin means and standard deviations of the errors.
-For the different datasets, we manually selected the bin width based on the range of predicted uncertainties.
-The exact bin widths can be found in Section S3.
+For the calibration curves and extended reliability diagrams, we calculate the uncertainty estimations and prediction errors on the candidate pool, with the respective GPR models trained on the 1000 training samples.
+In the extended reliability diagrams we separate the samples into equidistant bins by their uncertainty estimations and calculate the bin means and standard deviations of the errors.
+For the different datasets, we manually select the bin width based on the range of predicted uncertainties.
+The exact bin widths can be found in table S1 of the supplementary material.
 We compare the calculated bin means and standard deviation with the theoretical values of Eq.
-15.
+14.
 We only show the bin mean and standard deviation, if the number of samples inside one bin is greater than 50.
-For better visibility, we manually adjust the axis limits of the correlation plots, therefore a few scatter points are not displayed.
+For better visibility, we manually adjust the axis limits of the extended reliability diagrams, therefore a few scatter points are not displayed.
 For all uncertainty sampling tests, we use the fixed optimized hyperparameters and the test set, as described above.
-From all samples except the test samples we randomly choose 200 samples as initial training samples.
-3.1 rMD17 benzene
-3.1.1 Correlation between uncertainty estimate and prediction error
-(a) GPR with Coulomb
-(b) GPR with SOAP
-Figure 3: Correlation of uncertainty and error across different uncertainty estimations for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset.
-The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq.
-15.
+For the initial model, only 200 training samples are used, randomly selected from the dataset excluding the 2,000 test samples.
+3.1 Poor global calibration for ensemble uncertainties
+(a) Benzene
+(b) SMA
+Figure 3: Calibration curves of different uncertainty measures of GPR with SOAP for rMD17 benzene and WS22 SMA.
+Calibration curves plot the predicted proportion of test data expected to fall within $\alpha$-prediction intervals (horizontal axis) against the observed proportion of test data that actually falls within these intervals (vertical axis), iterating over different values of $\alpha$.
+In Fig.
+3, calibration curves for all three uncertainty measures are shown.
+We exemplary show calibration curves of GPR with SOAP for rMD17 benzene and WS22 SMA.
+The calibration curve for the GPR standard deviation is nearly diagonal and has a low miscalibration area, indicating relatively good calibration across the dataset.
+In contrast, the two sets and bootstrap uncertainty measures show much poorer calibration, with significantly higher miscalibration areas.
+In Fig.
+4 and in Figs.
+S3-7 of the supplementary material calibration curves for all uncertainty measures across all molecules are shown for GPR with Coulomb as well as GPR with SOAP.
+In nearly all cases, the GPR standard deviation demonstrates much better global calibration, with notably lower miscalibration areas compared to the other two uncertainty measures.
+The only exception is GPR with Coulomb for benzene.
+Here, the calibration curve for the GPR standard deviation shows a relatively high miscalibration area, which exceeds that of the two sets uncertainty measure.
+3.2 Local calibration analysis reveals model bias
+While the calibration curves indicate good global calibration for the GPR standard deviation, this does not mean that the model provides a well-calibrated uncertainty estimation for every prediction.
 Fig.
-3 shows correlation plots of different uncertainty estimators for benzene of the rMD17 dataset.
+4 presents calibration curves together with extended reliability diagrams for GPR with Coulomb across all molecules for the GPR standard deviation.
+As described above, the calibration curves indicate good global calibration for all molecules except benzene.
+However, when predictions are grouped by their uncertainty, the extended reliability diagrams reveal poor calibration for specific groups of predictions.
+While the standard deviation of errors often aligns rather closely with the theoretical value, the mean value of errors can deviate strongly from zero, particularly for predictions with high uncertainties.
+Assuming that the errors within separate bins, for which the means and standard deviations are calculated, are distributed somewhat close to normal, this can be interpreted as a systematic shift in the error distributions.
 In Fig.
-3(a), results for GPR with Coulomb representations are presented.
-For the GPR standard deviation and the bootstrap uncertainty, an increasing uncertainty correlates with a broader distribution of errors, as seen by the increasing bin standard deviations.
-Additionally, the bin means deviate further from zero with increasing uncertainty, indicating a bias in the model’s predictions.
-This suggests that both uncertainty estimators can identify samples for which the model has a high bias.
-The correlation between higher uncertainty and more broadly distributed errors as well as a higher bias means that for predictions with higher uncertainty higher errors are more likely.
-Still the observed error distributions do not align with what is predicted by the GPR models, making error ranges derived from these uncertainty estimations inaccurate.
-The systematical bias can not be capture by the uncertainty estimations.
-Besides that, for the GPR standard deviation, the bin standard deviations are smaller then the theoretical values, with an increasing deviation with increasing GPR standard deviation.
-Therefore, the errors are more narrowly distributed than predicted, which may partly compensate for the uncaptured bias when calculating confidence intervals.
-For the bootstrap uncertainty, for bins with lower uncertainty, the bin standard deviations exceed the theoretical values.
-Therefore the errors are distributed more broadly than predicted which will result in underconfident error intervals.
-For bins with higher uncertainty, the bin standard deviation aligns with the theoretical values, while the bin means deviate from zero, which will result in overconfident confidence intervals.
-Meanwhile, the two sets uncertainty shows errors consistently distributed around zero across all bins, with nearly identical standard deviations, providing no meaningful information about the error.
+S8 of the supplementary material, we explicitly show the error distributions within selected bins for GPR with Coulomb applied to Aspirin and O-HBDI.
+Indeed, these errors rather closely follow normal distributions hat are shifted relative to the predicted normal distribution.
+This systematic shift leads to much higher errors than predicted and indicates a systematic bias in the respective predictions.
+Although this bias is not captured quantitatively by the uncertainty estimation, it can still be observed that, in most cases, there is a clear correlation between increasing GPR standard deviation and increasing bias.
+This means the GPR standard deviation can identify predictions with a high bias.
+In ML theory, the error of a model is often divided into an irreducible part resulting from the noise in the training data, as well as a bias and a variance component.
+Hüllermeier and Waegeman[57] describe the GPR standard deviation as a meaningful indicator of the total model uncertainty.
+However, if we have a model that has a bias in the form of systematical shifts of the predictive distributions in certain areas of the input space, it is not possible for the standard deviation of the predictive distribution to quantify these shifts.
+In general, a bias is associated with incorrect model assumptions, making it challenging to quantify the resulting error using the flawed model itself.
+For ensemble-based uncertainties, Heid et al.[49] discussed that a bias cannot be quantified.
+Extended reliability diagrams for the GPR standard deviation of GPR with SOAP are shown in Fig.
+S5 of the supplementary material.
+In general, similar trends can be observed, although the bias is lower, and in some cases, the standard deviation of errors deviates more strongly from the theoretical value.
+Extended reliability diagrams for the two sets and bootstrap uncertainty of GPR wit Coulomb as well as GPR with SOAP can be found in Figs.
+S3-4 and S6-7 of the supplementary material.
+For the two sets uncertainty, it can be observed in all cases that errors are distributed with almost identical standard deviations and means across all bins.
+Consequently, the reliability diagrams reveal that the two set uncertainty not only shows poor global calibration but also provides no meaningful information to differentiate between predictions with different errors.
+A similar observation can be made for the bootstrap uncertainty in some of the cases.
+Figure 4: Calibration curves and extended reliability diagrams of the GPR standard deviation of GPR with Coulomb applied to different datasets.
+For the extended reliability diagrams, samples were separated into equidistant bins by their uncertainty, and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq.
+14.
+3.3 Uncertainty sampling results in worse error than random sampling
 In Fig.
-3(b), results for GPR with SOAP representations are shown.
-Again for the GPR standard deviation and the bootstrap uncertainty, an increasing uncertainty correlates with broader distributions of errors, therefore higher errors become more likely with increasing uncertainty.
-In contrast to GPR with Coulomb matrices, the bin means for both uncertainty estimators are close to zero and do not systematically increase with increasing uncertainty.
-Therefore, the GPR model has almost no systematical bias in the prediction.
-For the GPR standard deviation also the bin standard deviations are close to the theoretical value, therefore confidence intervals for predictions derived of the uncertainty would be rather accurate.
-For the bootstrap aggregation, the predicted error distributions for low uncertainties are slightly too broad, while error predictions of samples with higher uncertainties become too narrow with increasing discrepancy to the theoretical value for increasing uncertainty.
-This results in a less significant correlation between the uncertainty and a high likelihood for high errors than observed with the GPR standard deviation.
-Similar as for Coulomb matrices, the error distributions for the bins of the two set uncertainty do not differ for different uncertainties.
-Therefore, the two sets uncertainty again does not provide any meaningful information about the error.
-3.1.2 Active learning via uncertainty sampling
-Fig.
-4 shows uncertainty sampling results for benzene of the rMD17 dataset.
-In Fig.
-4(a), results for GPR with Coulomb representations are shown and in Fig.
-4(b) results for GPR with SOAP representations are shown.
-Adding samples based on maximum uncertainty estimations does not lead to any improvement in MAE, and in most cases, it also fails to improve the maximum absolute error and test variance.
-Only in the range of 900 to 1200 training samples for GPR with SOAP, the model with randomly added samples shows worse test variance and maximum absolute error compared to the model where samples with the highest uncertainty estimations were added.
-The correlation analysis shows how some of the uncertainty estimators provide relevant information about the error.
-E.g.
-when the GPR standard deviation is used as uncertainty estimator, a higher uncertainty correlates with a higher likelihood for larger errors.
-For GPR with SOAP, the observed error distributions actually align well with what is predicted with the GPR standard deviation.
-For GPR with Coulomb, there is a correlation between predictions with the highest GPR standard deviations and predictions that have the highest bias.
-Still, neither for GPR with Coulomb nor for GPR with SOAP uncertainty sampling with GPR standard deviation does result in higher data efficiency.
-In both cases, it can be seen that adding samples with the maximum absolute error improves all metrics significantly, compared to adding samples randomly.
-This shows that, in theory, there is a significant potential for a higher data efficiency.
-With the absolute error, or a potential uncertainty estimator which linearly correlates with the error, a prediction with a high error has a high uncertainty.
-In contrast, if the uncertainty estimator is the standard deviation of a predictive distribution, the errors of predictions with a high uncertainty follow a distribution in which high errors have a high likelihood but still also low errors are likely.
-This means many samples identified as maximally uncertain actually have a low error.
-By adding these samples, the model becomes certain about them, but it was predicting them with a low error already before.
-This might explain why no improved data efficiency is observed with the GPR standard deviation or bootstrap uncertainty as uncertainty estimator.
-(a) GPR with Coulomb
-(b) GPR with SOAP
-Figure 4: Uncertainty sampling results for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset.
-We randomly selected 200 samples as initial training samples and 2000 samples as test samples.
-From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample.
-Additionally, we randomly added a new sample in every iteration, and we added the sample with the highest actual absolute error.
-During each iteration, we compute the MAE, the maximum absolute error and the variance for the test set.
-3.2 Comparison for several molecules
-3.2.1 Correlation between uncertainty estimate and prediction error
-In Figure 5 correlation plots for the GPR standard deviation for different datasets are shown.
-In Figure 5(a), results for GPR with Coulomb are presented.
-In all cases, the bin means of bins with higher uncertainties deviate from zero, indicating a systematical bias of the respective predictions.
-Like it was observed for benzene, there is a correlation between increasing GPR standard deviation and increasing bias.
-In most cases, the bin standard deviation aligns rather well with the theoretical values.
-However, for SMA, it is observed that the bin standard deviations for bins with higher GPR standard deviations are smaller than the theoretical values.
-The deviation increases with increasing GPR standard deviation.
-The same can be observed to a smaller extent for O-HBDI.
-Therefore, the errors are more narrowly distributed than predicted.
-This may partly compensate for the uncaptured bias when calculating confidence intervals.
-Still, it can be observed that for predictions with high GPR standard deviation, the high bias dominates, which would lead to overconfident confidence intervals.
-In Fig.
-5(b), results for GPR with SOAP are presented.
-In all datasets, most of the bin means differ from zero, indicating that predictions with the GPR standard deviation of such bins have a bias.
-For SMA and O-HBDI, an increasing bias correlates with increasing uncertainty.
-For Aspirin, a clear correlation between increasing uncertainties and increasing bias cannot be observed.
-For SMA and Aspirin, the bin standard deviations align rather well with the theoretical values.
-Still, because of the uncaptured bias, resulting confidence intervals would be overconfident.
-For O-HBDI, the bin standard deviations for bins of low uncertainty align with the theoretical values, while for bins of high uncertainty the error distribution is much broader than predicted by the GPR model.
-This, combined with the uncaptured bias, would lead to highly overconfident confidence intervals.
-Overall, the correlation between higher uncertainties and more broadly distributed errors, as well as a higher bias in most cases, show that there is a clear correlation between uncertainty and a higher likelihood for higher errors.
-Still the bin means and bin standard deviations of the error distributions do often not align with the theoretical values which are predicted by the GPR model, therefore no accurate confidence intervals can be derived from the GPR standard deviation.
-Specifically, predictions with a high GPR standard deviation often have a high bias, which dominates the error and is not captured by the GPR standard deviation, which would lead to highly overconfident confidence intervals for such predictions.
-Here, we only compare correlation plots for the GPR standard deviation across different datasets, as it is the most widely used and discussed uncertainty estimator in the context of GPR based MLIPs.
-Results for two sets and bootstrap aggregation uncertainty estimations can be found in the supplementary material Section S4.
-In all cases, for the two sets uncertainty, the same picture as for benzene was observed: in all bins, the errors are distributed around zero with almost the same bin standard deviation.
-Therefore, again, the two sets uncertainty does not provide any meaningful information about the actual error.
-Similarly, correlation plots for the bootstrap aggregation uncertainty for Aspirin and SMA show very similar error distributions for all uncertainties, although a slight correlation between uncertainty and how much the bin means differ from zero, as well as how broad the error distributions are, can be observed.
-For O-HBDI, the bin standard deviations are larger than the theoretical values, with a correlation between increasing deviation and increasing uncertainties.
-The bin means of bins with large uncertainties also differ from zero.
-3.2.2 Active learning via uncertainty sampling
-In Fig.
-6, the MAE for every iteration of uncertainty sampling for different datasets is shown.
-For all cases, beside porphyrin with Coulomb matrices, it can be observed that uncertainty sampling with the GPR standard deviation results in a worse MAE than adding samples randomly.
-Among all cases, porphyrin with Coulomb matrices is the only one where uncertainty sampling results in a similar MAE than adding samples randomly.
-In contrast to the other cases, similar to benzene with SOAP, porphyrin showed an error distribution in the correlation plot that aligns well with the theory.
-The two sets uncertainty performs very similar to adding samples randomly, what was expected as the correlation plots show how they do not provide any meaningful information about the error.
-Uncertainty sampling with the bootstrap aggregation uncertainty leads to a worse MAE compared to random sampling in all cases beside SMA with SOAP.
-Results for the maximum absolute test error and the test variance can be found in Section S5.
-The maximum absolute error of the test set improves for uncertainty sampling with GPR standard deviation in all cases, besides for aspirin.
-The test variance in most cases is similar to random sampling or deteriorates, only for O-HBDI a slight improvement can be observed.
-When training samples with the highest absolute error were added, all three metrics improved significantly for all cases, showing that in principle there is quite some potential for an improved data efficiency.
-(a) GPR with Coulomb
-(b) GPR with SOAP
-Figure 5: Correlation of the GPR standard deviation and error across different datasets for GPR with Coulomb and GPR with SOAP.
-The samples were separated into equidistant bins by their uncertainty, and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq.
-15.
+5, the MAE for every iteration of uncertainty sampling across all molecules is shown.
+In most cases, uncertainty sampling with the GPR standard deviation results in a worse MAE compared to adding samples randomly.
+Only for benzene and porphyrin with Coulomb matrices, uncertainty sampling results in a similar MAE compared to adding samples randomly.
+Note that in the local calibration analysis, this were exactly the cases where the extended reliability diagrams did not show a significant bias for predictions with high uncertainties.
+The two sets uncertainty performs very similar to random sampling, as expected, since the extended reliability diagrams demonstrate that it does not provide any meaningful information about the error.
+Interestingly, uncertainty sampling with the bootstrap aggregation uncertainty results in a worse MAE than random sampling in almost all cases.
+When samples corresponding to predictions with the highest absolute error were added to the training set, the MAE improved significantly for all cases, showing that in principle there is a relevant potential for an improved data efficiency through better sampling strategies.
+Our findings align well with those of Uteva et al.[9], who reported that uncertainty sampling with the GPR variance performs worse than uncertainty sampling with the two sets uncertainty, adding samples corresponding to predictions with the highest absolute error, and training with maximum Latin hypercube data.
+The reduced data efficiency observed when sampling a fixed configuration space using uncertainty sampling suggests that, in on-the-fly active learning (AL), where uncertainty estimations guide the exploration of the configuration space, data efficiency may also be compromised.
+In contrast, Vandermause et al.[3] demonstrated increased data efficiency with their on-the-fly AL scheme applied to five-element high-entropy alloy structures.
+Even if the MAE is not improved by uncertainty sampling, a model trained via uncertainty sampling may perform more consistently across different data points in the test set.
+To analyze this, we present results for the maximum absolute test error and the test variance in Figs.
+S9 and S10 of the supplementary material.
+Indeed, the maximum absolute error of the test set improves for uncertainty sampling with the GPR standard deviation in all cases, except for aspirin.
+The test variance, however, is similar to that of random sampling in most cases or deteriorates, with improvements observed only for benzene and O-HBDI with Coulomb.
+When training samples corresponding to predictions with the highest absolute error were added, both metrics improved significantly for all cases.
 (a) GPR with Coulomb
@@
 During each iteration, we compute the MAE.
-3.3 Correlation analysis after applying active learning
-It may be surprising that, in most cases, uncertainty sampling with the GPR standard deviation leads to a significantly worse MAE on the test set compared to random sampling.
-In Figure 5, we observe that predictions with the highest GPR standard deviation are having the highest bias.
-This would suggest that adding these configurations as training samples should be beneficial.
-The most extreme case is GPR with Coulomb matrices for aspirin, where the MAE after 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty is worse than for the initial model.
-To understand this behavior we show in Fig.
-7 the correlation plots for Aspirin of a model trained via 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty estimation, as well as after iteratively adding 800 training configurations with the highest absolute error and adding 800 additional training configurations randomly.
-Similar to the model trained with 1000 random samples of Fig.
-5, the initial model, which is trained with 200 random training configurations, shows an increasing negative bias for predictions with increasing GPR standard deviation, after GPR standard deviation exceeds a certain threshold.
-For predictions below this threshold, there is a range of rather low GPR standard deviation where the predictions have a positive bias.
-After adding 800 configurations randomly, this general form is still present.
-After uncertainty sampling, on the other side, the bias of predictions with a high standard deviation vanishes, while the positive bias, as well as the standard deviation of the errors, of predictions with low standard deviation, increases.
-A possible explanation for this is that, if we add the configuration for which the model predicts the highest GPR standard deviation in every iteration, we add the configuration which is most distant to the current training set.
-After many iterations, we have a training set with configurations which are maximally distant to each other.
-The training set now covers the borders of the configuration space, for which the initial model was showing a high bias, very well.
-However, predictions for some areas of the configuration space, which are less distant to the initial training configurations, are getting even worse, compared to the initial model with much less training configurations, when the training set overemphasizes the edges.
-In contrast to training via uncertainty sampling and adding training samples randomly, iteratively adding configurations for which the model has the highest absolute error results in a model which has a low error for predictions over the full range of different GPR standard deviations.
-Figure 7: Correlation plots for GPR with Coulomb representations for Aspirin after 800 iterations of uncertainty sampling.
-The left plot in the first row shows the correlation plot with predicted GPR standard deviation vs.
-actual error for all configurations of the candidate pool of the initial GPR model that was trained with 200 random configurations.
-The right plot in the first row shows the uncertainty sampling results for 800 iterations.
-In the bottom row, correlation plots are shown for a GPR model trained with 800 iterations of uncertainty sampling with the GPR standard deviation as uncertainty estimation, as well as after iteratively adding 800 configurations with the highest absolute error and adding 800 additional training configurations randomly.
-4 Discussion
-The objective of this study was to evaluate different uncertainty estimators of GPR based MLIPs.
-To understand how uncertainty estimations correlate with the actual error, we plot both against each other and analyze if the errors are distributed according to the theoretical distributions, that result from the predictive distributions.
-For the GPR standard deviation, a clear correlation between an increasing uncertainty and increasing likelihood for larger errors is observed.
-For the two sets uncertainty, no correlation is observed, and the bootstrap uncertainty only shows a clear correlation in a few cases.
-However, even in cases where a correlation is observed, the actual error distributions often do not align with the error distributions that are predicted by the model.
-Therefore, the predictive distributions cannot quantify the error with valid confidence intervals.
-This also means it will not be possible to propagate uncertainty in a meaningful way.
-If the predictions of an MLIP are used to calculate further quantities, propagating uncertainty will not result in an accurate confidence intervals for the derived quantity.
-Still, if the uncertainty correlates with a higher likelihood for larger errors it provides a useful number, although a proper threshold might have to be found empirically.
-We found that specially predictions with higher uncertainties often will be highly overconfident, because the error is dominated by a high bias which is not captured quantitatively by the uncertainty.
-In contrast to our findings, Tran et al.38 found that the GPR standard deviation of their GPR model predicts underconfident confidence intervals for their exemplary dataset.
-The GAP model of Bartok et al.44 on the other hand predicts overconfident confidence intervals for their dataset of Ar dimers and trimers too, if the hyperparameters are optimized by maximizing the marginal likelihood, like we do.
-When optimizing the hyperparameters by optimizing the leave-one-out cross validation likelihood their GPR model predicts underconfident confidence intervals.
-In ML theory, the error of a model is often separated into an irreducible part which results from the noise in the training data, as well as a bias and a variance part.
-We observe that predictions with a high uncertainty often have a high bias, which is not captured by the uncertainty in a quantitative way.
-Hüllermeier and Waegeman54 describe the GPR standard deviation as meaningful indicator for the total model uncertainty.
-However, if we have a model that has a bias in the form of systematical shifts of the predictive distributions in some areas of the input space it is not possible, that the standard deviation of a predictive distribution quantifies this shifts.
-Therefore, in general, a bias is associated with incorrect model assumptions, which makes it challenging to quantify the resulting error with the flawed model itself.
-For ensemble based uncertainties Heid et al.47 discussed that a bias can not be quantified.
-When sampling a fixed configuration space with uncertainty sampling, and the GPR standard deviation is used as uncertainty estimator, the test MAE does not improve compared to random sampling in cases where the predictive distributions aligned well with the error distributions, while the maximum absolute error and the test variance sometimes improve compared to random sampling.
-For cases where the predictions with increasing uncertainty have an increasing bias, which are most of the cases, the MAE is getting worse compared to random sampling.
-This aligns well with the findings of Uteva et al.9, where uncertainty sampling with the GPR variance performs worse than uncertainty sampling with the two sets uncertainty and adding samples with the highest absolute error, as well as training with maximum Latin hypercube data.
-This behaviour might be explained by the fact that the GPR standard deviation finds a maximally distant training set, which results in a model that overemphasizes the border of the configuration space, that was originally sampled in the fixed dataset.
-The borders of the configuration space, which is represented in the fixed dataset might be sampled more sparse while other areas are sampled more densely.
-A model that overemphasizes the border might perform slightly worse on samples of other areas of the configuration space, which are more frequent.
-While this means that uncertainty sampling can result in an increased average model error inside the fixed configuration space, it is likely to be beneficial for extrapolation tasks.
-This observation was also made by Podryabinkin et al.7 when they trained an MLIP with the most uncertain configurations of a fixed dataset, according to the D-optimality criterium.
-Compared to random sampling, the model shows a slightly increased root mean squared error and a slightly decreased maximum error, when testing on a test dataset, which was previously selected randomly from the fixed dataset, compared to random sampling, while MD simulations with this model are much more stable.
-The reduced data efficiency, when sampling a fixed configuration space with uncertainty sampling, suggest that for on-the-fly AL, where uncertainty estimations guide the exploration of configuration space, data efficiency may also be compromised.
-Contrary to this, Vandermause et al.3 were showing an increased data efficiency with their on-the-fly AL scheme applied to 5-element high entropy alloy structures.
-Note that our tests were performed for energy predictions without forces.
-Other studies have reported that uncertainty estimations based on force predictions can be beneficial, because as forces are the derivative with respect to the atom coordinates, they carry more local information.
-Besides that, in general a higher model accuracy also leads to better uncertainty estimations.
-For the SOAP representations, parameters, which lead to higher resolution of the atomistic environments, would be possible.
-Still, we already use an atomistic kernel without sparsification, which in general leads to a much more accurate model than the classical GAP model with sparsification33.
-In cases where the error distributions align well with the predictive GPR distributions, we find that data efficiency does not improve.
-If error distributions align with the predictive GPR distributions, even the errors of predictions with a high uncertainty are distributed around zero and many configurations for which the model has a high uncertainty actually have a low error.
-When we add samples with the actual highest error, we observe a significantly improved data efficiency.
-Therefore, if we would have a more linear correlation, i.e.
-the error is directly proportional to the uncertainty, between uncertainty estimations and errors instead of a correlation between uncertainty and how broad error distributions are distributed around zero, we might archive a much better data efficiency.
-One approach to archive this is to average over atomistic uncertainties55.
-If we would have calculated the bin averages of the absolute error in our correlation plots instead of the error, for cases where the predictive GPR distributions aligned well with the error distributions we would see a linear correlation between uncertainty and absolute error.
-However, this linear correlation is only archived because we average over many samples, for a single sample the most likely absolute error is zero.
-For atomistic models, like GPR with SOAP, the energy prediction can be seen as a sum of atomistic energy predictions of atomistic submodels.
-If forces are considered the force prediction for every atom can be seen as an atomistic submodel too.
-Now averaging the uncertainty prediction of these atomistic submodels can lead to a linear correlation between uncertainty estimation and absolute error.
+3.4 More general model after uncertainty sampling?
+The significantly worse MAE for uncertainty sampling with the GPR standard deviation compared to random sampling might seem surprising, as we observe a correlation between increasing GPR standard deviation and increasing bias.
+This suggests that adding configurations corresponding to the highest GPR standard deviation as training samples should be beneficial.
+To analyze this behavior, in Fig.
+6 we show extended reliability diagrams after uncertainty sampling with the GPR standard deviation and after random sampling for GPR with Coulomb applied to Aspirin and SMA.
+For SMA, we show the plots for the models after all 4800 iterations, while for Aspirin, the plots are shown after only 800 iterations, because at that point, the model trained via uncertainty sampling with the GPR standard deviation performs even worse than the initial model.
+Additionally, we show the plot for the initial GPR model trained with 200 random samples.
+In both cases, it can be observed that the initial model not only has a high bias for predictions with high uncertainties but also a smaller bias in the opposite direction for predictions with very low uncertainty.
+After uncertainty sampling, the high bias for predictions with high uncertainty is eliminated in both cases.
+However, the bias for predictions with low uncertainty is higher than after random sampling and, for Aspirin, even higher than for the initial model.
+The standard deviation of errors of the respective bins with low uncertainty is higher than after random sampling too.
+These observations could be explained by the fact that uncertainty sampling with the GPR standard deviation finds a maximally distant training set.
+This may result in a model that strongly emphasizes the borders of the configuration space originally sampled in the fixed dataset.
+In such fixed datasets, the borders of the configuration space might be sampled more sparsely, while other areas are sampled more densely.
+A model that emphasizes the borders strongly might perform slightly worse on samples from other areas of the configuration space, which are more frequent.
+While this means that uncertainty sampling can result in an increased average model error within the fixed configuration space, it is likely to be beneficial for extrapolation tasks.
+This observation was also made by Podryabinkin et al.[7], who trained an MLIP using the most uncertain configurations of a fixed dataset, based on the D-optimality criterion.
+Compared to random sampling, the model shows a slightly increased root mean squared error and a slightly decreased maximum error when tested on a dataset randomly selected from the fixed dataset.
+However, MD simulations using this model were significantly more stable.
+(a) Aspirin
+(b) SMA
+Figure 6: Extended reliability diagrams of GPR with Coulomb representations for aspirin and SMA are presented.
+These plots compare the initial GPR models trained on 200 random configurations with those obtained after uncertainty sampling using the GPR standard deviation, as well as models updated by adding samples randomly.
+For better visibility, the plots for SMA use different bin sizes and axis limits between the initial model and those following uncertainty sampling.
+Note that the axis values also differ from those in Fig.
+4.
 4 Conclusion
-In this study, we evaluated how different uncertainty estimations for GPR based MLIPs relate to the actual error and if sampling a fixed configuration space by iteratively choosing configuration with the maximum uncertainty can improve the model accuracy compared to random sampling.
-We find that ensemble based uncertainties often do not provide meaningful information about the error, while for the GPR standard deviation a clear correlation between increasing uncertainty and increasing likelihood for higher errors is observed.
-Still also the GPR standard deviation for most datasets is not able to quantify the prediction error correctly, therefore no accurate confidence intervals for model predictions can be derived.
-While an increasing GPR standard deviation correlates with an increasing bias in the model predictions, this bias is not captured quantitatively by the uncertainty estimation and often dominates the error of predictions with high uncertainty, resulting in highly overconfident confidence intervals.
-When sampling a fixed configuration space with configurations for which a GPR model predicts maximum GPR standard deviation, training configurations are found which are maximally distant to each other.
-This may lead to a more general model, that performs better for extrapolation tasks.
-Still, such a model can overemphasize the borders of the configuration space of a fixed dataset and perform on average worse on the fixed dataset than a model based on random sampling.
+In this study, we evaluated calibration of different uncertainty measures for GPR based MLIPs and their performance when applied to active learning, where we iteratively choose the configuration with the maximum uncertainty from a fixed configuration space.
+Regarding calibration, it becomes evident how important it is to consider not only global calibration, which averages over the entire test dataset, but also local calibration, which considers specific groups of predictions.
+Ensemble-based uncertainty measures appear to be a poor choice for GPR-based MLIPs, as they already show poor global calibration.
+In contrast, the GPR standard deviation demonstrates relatively good global calibration.
+However, when calibration is evaluated locally, the GPR standard deviation can show very poor calibration.
+We find that, in particular, groups of predictions with a high GPR standard deviation often have a high bias that is not captured quantitatively by the uncertainty estimation.
+Therefore, without further knowledge, the GPR standard deviation should not be interpreted as a quantitative measure of potential error ranges, such as confidence intervals derived from the GPR standard deviation.
+Still, we find that an increasing GPR standard deviation correlates with an increasing bias and can therefore identify predictions with high bias and high error.
+When sampling a fixed configuration space by selecting configurations for which a GPR model predicts the maximum GPR standard deviation, the resulting training configurations are maximally distant from each other.
+This may lead to a more general model, that performs better in extrapolation tasks.
+However, such a model may overemphasize the borders of the configuration space in a fixed dataset, resulting in worse average performance on the fixed dataset compared to a model trained using random sampling.
 Acknowledgment
@@
 The authors would also like to acknowledge the support of the ‘Interdisciplinary Center for Machine Learning and Data Analytics (IZMD)’ at the University of Wuppertal.
+The authors also acknowledge the use of ChatGPT-4o, for assistance with language refinement, grammar corrections, and reformulations in the preparation of this manuscript.
+Data availability statement
+The data that support the findings of this study are openly available at the following URL: xxx
 S1 Hyperparameter optimization
-To optimize the GPR hyperparameters we maximize the marginal log likelihood (MLL) using gpytorch45.
+To optimize the GPR hyperparameters we maximize the marginal log likelihood (MLL) using GPyTorch[47].
 We use the ADAM optimizer with a learning rate of $lr=0.1$ and 200 steps.
-Since prior assumption of the hyperparameters, i.e.
-the hyperparameter starting values in the optimization process, can significantly influence the outcome, we perform cross validation for all combinations of possible starting values shown in Table S2.
-In all cases we use 1000 random samples to optimize the hyperparameters, for which we perform 5-fold cross validation with 5 repetitions.
-In every fold we optimize the hyperparameters with the respective training samples and starting values.
-Then we choose the hyperparamter starting values which resulted in the lowest loss on average over all repetitions and folds and use them as starting values for maximizing the MLL with all 1000 random samples.
+Since the initial guess of the hyperparameters in the optimization process can significantly influence the outcome, we perform cross validation for all combinations of possible initial guesses shown in Table S1.
+In all cases, we use 1000 random samples to optimize the hyperparameters, for which we perform 5-fold cross validation with 5 repetitions.
+In every fold, we optimize the hyperparameters with the respective training samples and starting values.
+Then we, choose the hyperparameter initial guesses, which resulted in the lowest loss on average over all repetitions and folds, and use them as intitial guesses for maximizing the MLL with all 1000 random samples.
 Hyperparameter Possible values
@@
 Noise $\sigma^{2}_{n}$ $10^{-4}$, $10^{-6}$, $10^{-8}$
-Table S1: Possible starting values for hyperparameters for the maximization of the MLL.
+Table S1: Possible initial guesses for hyperparameters for the maximization of the MLL.
 S2 Ensemble mean vs single model mean
@@
 (b) GPR with SOAP
-Figure S1: Correlation plots of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset.
-On the one hand the predictions are the mean of the submodel predictions of the two sets or bootstrap ensemble, on the other hand the predictions are from a single GPR model, trained on the full training set, from which subsets are used to train the ensemble submodels.
+Figure S1: Extended reliability diagrams of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset.
 (a) GPR with Coulomb
 (b) GPR with SOAP
-Figure S2: Correlation plots of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for aspirin from the rMD17 dataset.
-On the one hand the predictions are the mean of the submodel predictions of the two sets or bootstrap ensemble, on the other hand the predictions are from a single GPR model, trained on the full training set, from which subsets are used to train the ensemble submodels.
+Figure S2: Extended reliability diagrams of two sets and bootstrap aggregation uncertainty for GPR with Coulomb and SOAP representations for aspirin from the rMD17 dataset.
 S3 Bin widths
@@
 Table S2: Bin width of correlation plots for different molecules.
-S4 Comparison of Uncertainty error correlation between several molecules
+S4 Calibration analysis
+Figure S3: Calibration curves and extended reliability diagrams of the two sets uncertainty of GPR with Coulomb applied to different datasets.
+Figure S4: Calibration curves and extended reliability diagrams of the bootstrap uncertainty of GPR with Coulomb applied to different datasets.
+Figure S5: Calibration curves and extended reliability diagrams of the GPR standard deviation of GPR with SOAP applied to different datasets.
+Figure S6: Calibration curves and extended reliability diagrams of the two sets uncertainty of GPR with SOAP applied to different datasets.
+Figure S7: Calibration curves and extended reliability diagrams of the bootstrap uncertainty of GPR with SOAP applied to different datasets.
+S5 Explicit error distributions of separate bins
+(a) Aspirin
+(b) O-HBDI
+Figure S8: Predicted error distribution and distribution of true errors for selected bins of the reliability diagram of GPR with SOAP for Aspirin and O-HBDI.
+S6 Other metrics uncertainty sampling
 (a) GPR with Coulomb
 (b) GPR with SOAP
-Figure S3: Correlation of the two sets uncertainty and error across different datasets for GPR with Coulomb and GPR with SOAP.
-The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical error distributions.
-Note that for porphyrin and GPR with Coulomb the axis ranges and and bin width are adapted in comparison to the correlation plot for the GPR standard deviation for better visibility.
+Figure S9: Maximum absolut error of uncertainty sampling for GPR with Coulomb and SOAP representations for different datasets.
 (a) GPR with Coulomb
 (b) GPR with SOAP
-Figure S4: Correlation of the bootstrap uncertainty and error across different datasets for GPR with Coulomb and GPR with SOAP.
-The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical error distributions.
-Note that for porphyrin and GPR with Coulomb the axis ranges and and bin width are adapted in comparison to the correlation plot for the GPR standard deviation for better visibility.
-S5 Comparison of uncertainty sampling between several molecules
-(a) GPR with Coulomb
-(b) GPR with SOAP
-Figure S5: Maximum absolute error of every uncertainty sampling iteration for GPR with Coulomb and SOAP representations for different datasets.
-We randomly selected 200 samples as initial training samples and 2000 samples as test samples.
-From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample.
-Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error.
-(a) GPR with Coulomb
-(b) GPR with SOAP
-Figure S6: Test variance of every uncertainty sampling iteration for GPR with Coulomb and SOAP representations for different datasets.
-We randomly selected 200 samples as initial training samples and 2000 samples as test samples.
-From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample.
-Additionally we randomly added a new sample in every iteration and we added the sample with the highest actual absolute error.
+Figure S10: Test variance of uncertainty sampling for GPR with Coulomb and SOAP representations for different datasets.